### v1

# The building blocks of software work explain coding careers and language popularity

Abstract
Recent waves of technological transformation have fueled debates about the changing nature of work. Yet to understand the future of work, we need to know more about what people actually do in their jobs, going beyond educational credentials or job descriptions. Here we analyze work in the global software industry using tens of millions of Question and Answer posts on Stack Overflow to create a fine-grained taxonomy of software tasks, the elementary building blocks of software development work. These tasks predict salaries and job requirements in real-world job ads. We also observe how individuals learn within tasks and diversify into new tasks. Tasks that people acquire tend to be related to their old ones, but of lower value, suggesting that they are easier. An exception is users of Python, an increasingly popular programming language known for its versatility. Python users enter tasks that tend to be higher-value, providing an explanation for the language’s growing popularity based on the tasks Python enables its users to perform. In general, these insights demonstrate the value of task taxonomies extracted at scale from large datasets: they offer high resolution and near real-time descriptions of changing labor markets. In the case of software tasks, they map such changes for jobs at the forefront of a digitizing global economy.
Classification: Social Sciences/Economics
Keywords: human capital; tasks; software; economic complexity; networks
1 Introduction
Knowledge of the tasks that workers perform is central to understanding wages, productivity, and overall labor market dynamics (Acemoglu and Autor, 2011). Recent work has transformed our ability to do so by using complex network analysis — moving away from descriptions that rely on years of schooling, credentials, and work experience to the granular building blocks of human capital: skills and tasks (Anderson, 2017; Alabdulkareem et al., 2018; Hosseinioun et al., 2025). These studies focus on jobs, describing them as networks of skills or tasks using data on skill requirements and task descriptions extracted from large-scale surveys (one, 2025), expert assessments (Commission, 2017), job ads (Börner et al., 2018) or online freelance platforms(Anderson, 2017; Stephany and Teutloff, 2024). Yet, job descriptions tell us something about the nature of vacancies, not of the people who fulfill them. This limits our ability to understand how workers specialize and learn new things, processes critical to a successful career.
Focusing on the software sector, we propose a new approach that uses data from Stack Overflow (SO), a very large, online question-answer database. The core insight we leverage is that questions on SO describe coding problems that programmers encounter in their work. As such, they can be understood as instantiations of software-development job-tasks. By generalizing these descriptions of specific problems encountered at work, we can construct a taxonomy of canonical software tasks that is highly granular, flexible and based in data. Moreover, we can use this taxonomy to describe the human capital endowments of people, not the human capital requirements of jobs. That is, we ask what programmers do at work, not what they are supposed to do according to their job descriptions. This allows us to analyze how programmers learn new tasks and what the value of these tasks is. Moreover, programmers can choose from a wide menu of programming languages and we show that programming languages differ in the tasks they support. However, the mapping of tasks to programming languages changes over time, as illustrated by the rapid rise of Python as a general purpose language. Our task network shows that this rise exploits a remarkable strength of Python: it allows programmers to more quickly transition to the most valuable programming tasks.
As wages and careers have come to depend on how workers specialize (Hosseinioun et al., 2025), interact (Deming, 2017), complement colleagues (Neffke, 2019), and move across places and jobs (Frank et al., 2024), fine-grained taxonomies of tasks are increasingly valuable. Such taxonomies clarify how workers transition between occupations (del Rio-Chanona et al., 2021), how skills can substitute for or complement each other (Anderson et al., 2012; Neffke, 2019), and how occupations evolve over time (Nedelkoska et al., 2021).
In this paper, we focus on jobs in the global software development sector. Software production is a fast-changing and critical domain of the digital economy (Brynjolfsson and McAfee, 2014; Wachs et al., 2022; Juhász et al., 2024). Of the seven largest companies by market capitalization worldwide, three (Microsoft, Alphabet, and Meta) focus almost exclusively on software, while three others (Nvidia, Apple, and Amazon) derive substantial revenue from it. Demand for programming skills continues to surge: one in 11 U.S. occupations now requires coding (one, 2025), and software-related roles such as data scientist, information security analyst, and computer research scientist rank among the fastest-growing occupations (U.S. Bureau of Labor Statistics, 2025). As new hardware, programming languages, and specialized libraries reshape computing, fresh roles (e.g., web development, DevOps, data science, artificial intelligence) keep emerging (Rock, 2019; Papoutsoglou et al., 2019; Hemon et al., 2020). This changing landscape has sparked intense global competition for talent, prompting many major economies to designate software development as a strategic priority (Draghi, 2024; Kazim et al., 2021; Singh, 2024).
However, we currently lack a detailed understanding of the fast-changing human capital that is employed in this sector. One reason is that databases linking occupations to tasks (e.g., O*NET (one, 2025)) are often curated by human experts, making them relatively coarse and slow to update. In O*NET, for example, software-related tasks often focus on broad domains of application (e.g., “geoinformatics” or “biostatistics”), team dynamics (e.g., managing or collaborating with software engineers), or acquiring software products. Alternative data sources such as job advertisements or resumes are also challenging because job ads describe job requirements of ideal-type workers, not human capital endowments of actual workers, resumes can be aspirational, and both often focus on limited geographies. Likewise, platforms for online freelancing provide insights into how complex skill interdependencies affect wage premiums (Anderson, 2017; Stephany and Teutloff, 2024), but the extent to which these findings generalize to the broader labor market remains uncertain.
We address these gaps by analyzing millions of posts from SO—the largest question-and-answer forum for programmers (Anderson et al., 2012). Each post describes a specific technical challenge and includes tags that indicate the tools, languages, or concepts involved (Barua et al., 2014). These challenges reflect real-world coding tasks and users use their posting histories in part to signal their skills to prospective employers (Xu et al., 2020). Interpreting tags as a structured, albeit partial, description of posted questions, we aggregate questions to create a software-task taxonomy by identifying clusters of frequently co-occurring tags. Using this taxonomy, we examine how software tasks evolve in the labor market. We find that task value estimates, derived from survey data, significantly predict salaries in real-world job postings. Notably, wages vary greatly among these tasks, with higher-paying ones concentrated in more specialized or in-demand fields such as AI and machine learning. We also introduce a notion of task relatedness based on co-occurrence patterns in users’ posting histories. This measure of relatedness derived from interactions with the SO platform predicts similar co-occurrence trends in job ads, suggesting that our taxonomy captures a meaningful structure of the software development landscape.
At the individual user level, we observe two learning patterns. First, developers improve in the tasks they gain experience in, as measured by community feedback on their posts. Second, developers enter new tasks over their SO careers, and task relatedness strongly influences these entries: developers are fifteen times more likely to move into tasks related to ones they already know. At the same time, we find that developers usually add tasks with lower market value, which are presumably easier to master.
Finally we demonstrate that our fine-grained software task taxonomy sheds light on the evolution of software development as a whole. We observe that programming languages differ in the number of tasks for which they are used, following a nested pattern of tasks in programming languages reminiscent of similar patterns observed in ecology (Bascompte et al., 2003), economic development (Hidalgo and Hausmann, 2009; Tacchella et al., 2012; Mariani et al., 2019) and skill hierarchies (Hosseinioun et al., 2025). Accordingly, instead of specialization, where different languages are used to perform different types of tasks, we observe that tasks have an implicit hierarchy such that the most ubiquitous tasks can be performed in all languages, whereas rare tasks are only performed by the most versatile ones. A prime example of the latter is the Python programming language, a popular language that has seen significant growth in recent years (Economist, 2018). We show that this growth has been accompanied by changes in the way Python is used. First, Python has become increasingly versatile. From 2015, the number of tasks in which Python dominates as the preferred language for most programmers has rapidly grown, rendering Python into one of today’s premier general purpose languages. Second, Python enables qualitatively different paths for users to learn new tasks. Specifically, unlike users of most other languages, Python users often branch into more, not less valuable tasks. This observation could help explain Python’s success, despite its not being the oldest or fastest language: its ecosystem may make it easier for developers to tackle higher-value tasks.
2 Results
2.1 Constructing a space of software tasks
Software development tasks can be viewed as recurring categories of programming problems developers encounter. SO, with its vast repository of user-generated questions, offers a unique window into these real-world challenges. To organize this repository and facilitate retrieval, each question is labeled with one or more of 66,000 curated tags that indicate the specific tools, frameworks, or concepts involved.
We use these tag–question relationships to group related issues into “canonical” software tasks, applying a bipartite stochastic block model (SBM) (Peixoto, 2017, 2014; Gerlach et al., 2018) to detect communities in the graph of questions and tags (Fig. 1a). We ignore tags that refer to programming languages and use them later on to study how tasks are distributed across languages. To generate concise names for each task community, we prompt ChatGPT-4.0 for labels that capture the essence of the tags comprising each group (see Appendix B). This process yields a set of 237 software tasks, such as Develop NLP models and tools for text analysis, Establish secure server access and control, and Deploy and secure a scalable web app.
Figure 1: Mapping software tasks. a. Stylized depiction of the bipartite question-tag network. SBM groups tags into communities (tasks) that connect to similar sets of questions. ChatGPT-4.0 finds a common label that summarizes each community’s tag information. b. Task space. Pointwise mutual information ($\mathrm{PMI}$) expresses how surprisingly often two tasks are performed by the same users. UMAP embeds the resulting co-occurrence network in a 2-dimensional plane (the task space). c. Close-up on Develop AI models task, depicting the original network structure among the 10 most closely related tasks. d. Task values. Nodes are colored according to their task value, estimated from salary information in the SO 2023 developers survey. Darker shades indicate more valuable tasks. e. Table of the five most and least valuable tasks. For details, see Methods.
Because certain tasks complement each other, they often co-occur within the same job roles (Anderson, 2017; Alabdulkareem et al., 2018; Neffke, 2019; Stephany and Teutloff, 2024). To investigate this task bundling, we identify which tasks are frequently performed by the same SO users. This yields a network that links tasks based on their co-occurrence in user activity. To reduce dimensionality and facilitate visualization, we then embed this network in a two-dimensional plane.
The resulting task space contains several clusters of related tasks, each shown in a different color in Fig. 1b (see Methods). Many of these clusters group tasks with comparable labor market returns. This is illustrated in Fig. 1d, where, we plot task values inferred from self-reported wages in a large-scale survey run among SO users (see Methods). On the right side of the task space, for example, three clusters revolve around website development, typically linked to lower wages. By contrast, tasks in the top center—focusing on advanced iOS and Android development—offer significantly higher earnings. Additional high-value clusters include those for AI and machine learning (AI/ML), Advanced Programming Concepts, Cloud Computing, and DevOps. The table of Fig. 1e lists the five highest-paying tasks, three of which are related to iOS app development, and five lowest-paying tasks, all associated with more basic web development.
Notably, clusters that are conceptually linked—such as AI/ML and Statistics and Data Analysis, and iOS and Android development—tend to lie near each other in the task space. To investigate the network structure of the task space, we can examine individual tasks in greater detail. For example, focusing on the ten tasks most closely related to developing AI models, Fig. 1c reveals strong connections to other statistical tasks, as well as tasks involving image and natural language processing, data visualization, and packaging models into standalone applications.
2.2 Validation using job ads
How well do these software tasks capture real-world development work beyond SO? To find out, we examine approximately 50,000 software job ads from HackerNews [^1] https://www.hackernews.com/ (see Fig. 2a and Methods). Using an LLM, we parse each job ad’s text to extract salaries and the list of tasks required of successful candidates.
Next, we convert each job’s extracted tasks into a 237-dimensional binary vector, where each dimension corresponds to one of the SO tasks identified previously. We accomplish this by embedding both the job’s task descriptions and the SO task labels into a 384-dimensional space using sBERT(Reimers and Gurevych, 2019). We then match each extracted task to its closest SO task by comparing cosine similarities. In this way, each job ad becomes a vector of job requirements, expressed in terms of SO tasks. Finally, we combine this vector with task-value information and measures of task relatedness. Figure 2a (see Methods) illustrates this process.
Figure 2: Job ads. a. Schematic representation of the workflow to extract salary and task requirements from online job ads by prompting ChatGPT. Task requirements are converted to the 237-dimensional SO task vectors of Fig. 1 based on cosine similarities between text embeddings of task requirements and SO tasks. b. Prediction of task requirements. 40% of task vector elements are masked and grouped into 10 equally sized bins based on the fit of the masked task to the (unmasked) task requirements of a job (see Methods). The plot displays the estimated probability that tasks in a bin are required in the job. c. Predictions of wage offers. Jobs are grouped into equally sized bins based on the average value of required tasks. The vertical axis shows the estimated mean of the advertised wage offers. Vertical bars in panels b and c represent 95% confidence intervals.
The task space strongly predicts which tasks appear together in job ads. To demonstrate this, we combine all job-task vectors into a single dataset and randomly mask 40% of the entries. For each masked task, we then measure its relatedness to the remaining unmasked tasks for that job. As shown in Fig. 2b, the probability that a masked task is indeed required (i.e., has a “1” in the task vector) increases sharply with its relatedness to the other tasks, rising 70-fold from below 1% for unrelated tasks to over 12% for strongly related ones.
The tasks we identify also predict wage offers (Fig. 2c). Salaries in job ads increase alongside the average value of the tasks a job requires, and regression analysis confirms this: a 10% rise in the average value of these tasks is associated with a 9% increase in the offered wage (see Appendix C.2).
2.3 Task dynamics: growth, diversification, feedback
Figure 3: Task dynamics. a. Task user-share change from 2009 to 2022. Purple markers signal increases, orange markers decreases, in user-shares between 2009 and 2022. Marker transparency reflects the size of shifts: darker tones indicate larger changes in user shares. b. Estimated probability of diversifying into new tasks at different values of density, users’ relatedness-weighted experience in other tasks: $d_{\theta,u}=\sum_{\kappa}\frac{\mathrm{R}_{\theta,\kappa}}{\sum_{\tau}\mathrm{R}_{\theta,\tau}}X_{\kappa,u}$, where $X_{\kappa,u}$ denotes user $u$’s experience in task $\kappa$ and $\mathrm{R}_{\theta,\kappa}$ the relatedness between tasks $\theta$ and $\kappa$. c. Regression analysis of answer popularity on task experience. Popularity is measured as the number of votes the answer receives, $\log(\#\text{votes}_{a}+1)$ or whether or not the answer is the top answer to the question, task experience as the number of prior answers the user provided to questions on task $\theta$ in the preceding two years, $\log(\#\text{answers}_{u(a),\theta,t(a)})$. Control variables include $\log(\#\text{answers}_{q(a)})$ and $\log\left(\sum_{\alpha\in A_{q}}\#\text{votes}_{\alpha}+1\right)$, the total number of answers provided to question $q(a))$, and the sum of all votes across these answers. To avoid problems due to $\log 0$ values, we add 1 to counts that can evaluate to 0. The plot shows point estimates with their 95% confidence intervals. Purple markers refer regression analyses of whether or not an answer is the top answer, orange markers to analyses of the number of votes. d. Regression of whether or not a user will adopt a new task $\theta$, on the task value, $V_{\theta}$, controlling for the user’s density around the task, $d_{\theta,u}$. Purple markers refer to Python-related questions only, orange markers to the full sample.
Figure 3a illustrates how SO activity shifted across software tasks between 2009 and 2022. Notable growth appears in Web Frameworks, certain Android-related tasks, Cloud Computing, and AI/ML. Conversely, tasks associated with Advanced Programming Concepts (e.g., Develop a data structure library for efficient data manipulation, Develop a multimedia application in Windows, or Develop a thread-safe application to prevent race conditions), as well as Networking, Enterprise and Web App Development, have seen their relative importance on SO decline.
We can also apply task vectors to study changes at the individual level, studying how individual SO users develop expertise over time. To show this, we generate a task vector for each user by counting how many answers they contribute in a two-year interval for each of the 237 defined tasks (see Methods). This approach captures the intensity of engagement in different areas of software work. We use these user task vectors to examine how developers add new skills to their profiles over time. In particular, we test whether task relatedness—reflected in the task space—predicts which new tasks a user will adopt on SO. As shown in Fig. 3b, users are far more likely to pick up tasks closely related to those they already know, with the probability of transitioning to a “new” task rising from 0.9% (when the task is unrelated) to 14.3% (when it is highly related).
We also expect that a user’s task-specific experience should correlate with the popularity of their answers. Testing this hypothesis is complicated, since users often decide to answer a question only after gauging existing solutions—a process that introduces selection bias. In Appendix D.1, we address these biases using instrumental variables. Here, we mitigate them partially by focusing on only the earliest answer posted to each question. We then ask whether users with greater experience in tasks connected to the question receive more votes. Concretely, we regress a binary outcome—whether the answer is the highest-voted—on the user’s prior task experience, controlling for the question’s total number of answers and sum of votes. As a robustness check, we also predict the total votes an answer receives.
Figure 3c reveals a statistically significant, positive link between users’ task experience and the popularity of their answers. The estimates indicate that a 10% rise in task experience boosts the likelihood of an answer becoming the top-voted solution by 0.1 pp (±$\pm 0.00004$) and increases the number of votes it receives by 0.14% (±$\pm 0.006\%$). In the SI, section D.2 we show that when with sample-selection corrections, the estimated effects grow to 0.006 pp (±$\pm 0.0035$) and 0.9% (±$\pm 0.41\%$), respectively.
Users also need motivation to adopt new tasks, whether for higher potential rewards or because the tasks align closely with their existing skills. To examine these motivations, we regress task entry on the estimated task value (from the aforementioned survey) and the user’s density around the task, i.e., the relatedness-weighted average experience in other tasks the user engages with (see Methods). As shown in Fig. 3d, developers do not universally pursue high-value tasks—likely due to the higher technological barriers or learning costs associated with them—and tend to choose tasks that are more closely related to their current skill sets (see Appendix D.3). However, Python users follow a different pattern: task value has a positive effect on their decision to adopt a new skill, suggesting that Python’s versatility allows them to more readily acquire profitable tasks. This highlights Python’s distinct role in the software ecosystem: as a general-purpose language, it enables users to branch into new, higher-value tasks more easily, altering and expanding their career trajectories.
2.4 Tasks and programming languages
Programmers can draw on a wide variety of programming languages, each with different levels of suitability and popularity for specific software tasks. Our observation that Python users often diversify into higher-value tasks suggests that language choice significantly influences developers’ career paths. Yet debates persist about which languages best serve which tasks. Our task constructs promise a novel way to study language usage. To explore this, we use tags that refer to specific programming languages to associate SO posts with the languages they refer to (see Methods). This allows us to track which tasks a language supports.
Figure 4a presents a matrix illustrating which programming languages (used by at least 10 SO users) are used in each of the 237 SO tasks. The matrix’s triangular shape signals task nestedness, with an NODF value of 81.5 (Payrató-Borràs et al., 2020), indicating that tasks can be roughly ranked from general to specialized based on how widespread they are across languages. Moreover, although different languages may be developed with specific use cases in mind, in practice, this has not led to a strict specialization where tasks are divided between dedicated programming languages. Instead, the most specialized tasks are typically handled by languages that cover almost the full spectrum of tasks, while languages used for only a few tasks are largely confined to the most ubiquitous ones.
Figure 4: Programming languages. a. Task-language matrix. Elements are colored when at least 10 SO users have at least one answer post in the task-language combination. b. Number of tasks in which a programming language ranks as the top language in terms of SO users. The graph shows time-series for the largest six languages in terms of cumulative SO posts between August 2008 and June 2023. c. Python’s task footprint. Nodes are colored if Python ranks among the top 3 programming languages for the associated task. Fully colored: 1st rank, 50% transparency: 2nd rank, 75% transparency: 3rd rank.
Figure 4b illustrates how different programming languages compete for tasks over time. Focusing on the six largest languages on Stack Overflow, it tracks how many tasks each language dominates by user counts. The most prominent trend is Python’s steady climb: from just nine tasks in 2009, it rises to become the top language for over 80 of the 237 tasks by 2022, surpassing well-established options like C# and JavaScript. In the SI, section F, we further demonstrate how languages compete over tasks by focusing on the abrupt shift in iOS programming from Objective-C to Swift.
Figure 4c visualizes Python’s growth, highlighting all tasks in which it ranks among the top three languages. Initially focused on data science, Python spread between 2014 and 2018 into tasks ranging from app development to web-related work, DevOps, and cloud computing. Today, it is the dominant language for at least one task in every cluster. As detailed in Appendix E, this pattern is preserved when we re-weight languages by their GitHub script counts, a proxy for real-world usage—indicating that Python’s reach extends well beyond Stack Overflow.
3 Discussion
The growing complexity of the labor market has made measuring worker capabilities both more difficult and more important (Anderson, 2017). Here we demonstrated that data on individual problem solving can be used to generate a useful fine-grained taxonomy of tasks people do in software-related work, a key industry in the digital economy. These tasks or building blocks of software work can be given coherent economic values and capture co-occurrence relationships reflected in real world job advertisements. We use the tasks to observe individual learning, both within tasks and entrance into new ones. Moreover, the fine-grained detail of the task space allows us to explain an important macro-system outcome: the rise of Python as the leading all-purpose programming language. Our explanation for Python’s success is that it enables people to enter new tasks which have higher economic value. Ultimately, our results highlight how deepening our understanding of the microstructure of tasks offers new ways to interpret macro-level phenomena such as technology diffusion, wage dispersion, and the evolution of industry-wide practices.
A core insight from our results is that tasks exist in a space of related capabilities, echoing network-based accounts of human capital (Anderson, 2017). On the one hand, tasks are well-defined as distinctive categories of specific kinds of work, as evidenced by their distinct labor market values and that effective learning happens by practice within tasks (Arrow, 1962). On the other hand, tasks also have meaningful relationships with each other, evidenced by their non-random co-occurrence within job ads and observed patterns of individuals learning new tasks. The resulting networked taxonomy of software tasks demonstrates its value by providing a micro-to-macro explanation for the rise of the Python programming language: Python enables users to enter more valuable tasks.
Our findings connect with the literature on skill complexity, labor market transitions, and the organization of work, complementing studies that have emphasized how granular task analysis can reveal hidden dynamics in career mobility, wage disparities, and skill complementarities (Autor et al., 2003; Anderson, 2017; Alabdulkareem et al., 2018; del Rio-Chanona et al., 2021). Previous research however, tends to use standardized task classifications such as O*NET to define tasks and applies them to curated job descriptions, limiting the granularity and responsiveness of task taxonomies. In contrast, by extracting tasks directly from real-world problem-solving interactions on Stack Overflow, our approach captures the immediate and practical realities of skill application at a global scale. In rapidly changing sectors such as software engineering, tasks derived from activities can explain surprising developments like the rise of the Python programming language as a technology which enables diversification into more valuable tasks. This complements existing work on task-relatedness and collective learning.
We highlight three possible applications of our method and taxonomy. First we have demonstrated that microdata on activity can be fruitfully aggregated into tasks to describe skill acquisition, development and transitions. Tasks from any sector in which large-scale fine-grained activity on individual behavior at work is available could be categorized in a similar manner. Second, employers in such sectors can use our approach to identify skills related to tasks their employees actually carry out, enabling better labor market matching outcomes or internal skill development programs. Finally, educators can better understand the labor market landscape of the software industry using our taxonomy.
There are several limitations to our study, which may be tackled in future work. First, we have focused exclusively on programming-related tasks. However, software developers need to be able to carry out many other tasks, such as managing teams or reporting results. Although SO contains some postings about organizational aspects of software development, its emphasis is on technical subjects. To explore how programming tasks are combined with other types of tasks in software jobs, future studies could use job ads to compare the co-occurrence of our programming tasks with tasks related to different kinds of soft skills.
Second, although the question-answer structure makes SO ideal for identifying software tasks, it may be less suited for descriptions of macro-level trends and individual career paths. In Appendix E, we compare user numbers on SO to GitHub. This analysis suggests that in terms of capturing software developers, both datasets are closely aligned. However, it is impossible to judge the representativity of SO of overall software activity at the level of tasks, especially as Stack Overflow is thought to be declining (del Rio-Chanona et al., 2024). In Appendix E, we therefore rescale the activity in different programming languages such that they reflect their presence on GitHub. However, in future work, we plan to define software tasks directly in GitHub.
Third, we opted to create a task taxonomy that is time-invariant. However, an analysis of how job tasks change in software development would provide a more vivid picture of the software industry. Such an analysis could also expand the size of our job ad dataset. Although sufficient for the validation exercise we conducted, a larger dataset would allow us to see how tasks in software development and their value change over time.
Notwithstanding these limitations, our study demonstrates how question-and-answer data can inform us about the evolution of work by quantifying the abstract notion of a task. The application to the software sector sheds light on novel aspects of high skill labor markets at high granularity, at scale, and in near real-time. Moreover, programming tasks are rapidly diffusing beyond pure software development firms and understanding these tasks and their evolution is pivotal to understanding the future of work in an increasingly digital economy.
4 Methods
4.1 Data
Stack Overflow Posts:
Our main dataset is the corpus of questions and answers (together “posts”) on Stack Overflow from the site’s launch in August 2008 to June 2023. The 23 million questions in this dataset are labeled by tags from a curated system of about 66,000 tags referring to software concepts and technologies such as key-value, scipy and svm. For each tag, SO provides a textual description, typically consisting of a few sentences. Tags are heavily moderated by humans and automated processes, and are deduplicated through a dictionary of synonyms. We focus on tags that are used at least 1,000 times. We split tags into a two groups: 5,083 general tags and 282 tags that refer to programming languages, defined broadly. For the latter, we merge different versions of the same programming language, converting, for instance python-3.x into the generic tag python. When we define tasks, we use all available posts, with answers inheriting the tags of the questions they answer. In all other analyses, we only consider answers from users with a total of at least 10 answer posts.
Stack Overflow Develop Survey:
Stack Overflow provides access to yearly surveys of its users, the Stack Overflow Developer Survey, with anonymous information on salaries, as well as on tools and languages used that can often be linked to tags in SO. We use the survey data from year 2023, which contains around 12,000 participants who report their salaries to estimate task values.
HackerNews Job Ads:
We scraped job advertisements from “hnhiring.com”, an index of jobs from the website Hacker News’ “Who is Hiring?” posts. This yielded a total of 50,998 job advertisements. We use approximately 46,000 ads to study task co-occurrence and 3,949 jobs to test the ability of tasks to predict salaries.
4.2 Building software tasks
4.2.1 Identifying programming languages
To identify tags that refer to programming languages, we rely on wiki pages, which are available for most common tags. We first filter the pages to find tags that include “language” or “framework” among their keywords. We then manually review all the tags to select those that refer to programming languages, yielding a total of 282 programming languages.
Some programming tags refer to specific versions of a language, such as “python-3.x” or “python-3.7”. When calculating the nestedness of the task-language matrix, we merge all tags referring to different versions of the same programming language. Yet, there is some ambiguity to what counts as a programming language, such as frameworks and developer environments. We remove such tags when calculating nestedness, basing this estimate on a task-language matrix with 114 programming languages.
4.2.2 Identifying software tasks
To identify canonical software tasks, we express the relation between questions and their tags as a bipartite network with 18,154,593 question nodes that connect to 5,083 tags through 33,378,590 edges. We then apply a bipartite Stochastic Block Model (SBM) to identify communities of tags.
SBMs essentially estimate a coarsened version of a network. To do so, they identify communities of nodes and parameters that represent parsimonious descriptions of edge distributions within and between communities. To sharpen the boundaries of inferred communities, we remove $20\%$ of tags that have the weakest own-community attachment (see Appendix section A). Dropping communities with fewer than 3 tags, we retain 237 tasks, connected to 4,054 tags, out of originally 247 detected communities.
4.3 User task vectors
We use the tasks extracted from the question-tag network to describe the activity of individual SO users as a 237-dimensional vector. Entries of this vector count how many answers the user has provided to questions that are labeled with tags that are associated with each corresponding task. Specifically, we express the user’s task experience in time period $t$ as:
$\vec{T}_{u}^{t}=\begin{bmatrix}X_{u,1}^{t}\\ ...\\ X_{u,\theta}^{t}\\ ...\\ X_{u,N_{\theta}}^{t}\\ \end{bmatrix},$ (1)
where $N_{\theta}=237$ the total number of tasks and $X_{u,\theta}^{t}$ is the number of answers posted in period $t$ by user $u$ to questions labeled with a tag belonging to the community of tags associated with task $\theta$.
4.4 Constructing a Task Space
To calculate the relatedness between tasks, we analyze which tasks are often associated with the same users. To do so, we collect the task vectors of all users in an $(N_{\theta}\times N_{u})$ matrix T, with $N_{u}$ the number of users in the dataset with at least 10 answers and $N_{\theta}=237$ the total number of tasks. Next, we ask how often two tasks co-occur within the same users by multiplying $(\textbf{T}>0)$ with its transpose: $(\textbf{T}>0)(\textbf{T}>0)^{\prime}$. Finally, we calculate the Pointwise Mutual Information ($\mathrm{PMI}$) of probability $p_{\theta,\kappa}$ that task $\theta$ and $\kappa$ co-occur in the same user to capture the amount of (information-theoretic) surprise involved in observing the joint probability $p_{\theta,\kappa}$ against a benchmark where draws of tasks $\theta$ and $\kappa$ are independent:
$\mathrm{PMI}_{\theta,\kappa}=\log\frac{p_{\theta,\kappa}}{p_{\theta}p_{\kappa}},$ (2)
where $p_{\theta}$ and $p_{\kappa}$ are marginal probabilities, i.e., $p_{\theta}=\sum_{\gamma}p_{\theta,\gamma}$ and $p_{\kappa}=\sum_{\mu}p_{\mu,\kappa}$.
To estimate $\mathrm{PMI}$ values and their credible intervals, we apply a Bayesian statistical framework developed by (van Dam et al., 2023).
The task space is based on a weighted network where nodes represent tasks and edges the relatedness between them. Because our focus is on relatedness as expressed in a surprisingly large observed probability that two tasks co-occur, we ignore tasks that co-occur less frequently than our random null model predicts by setting all negative $\mathrm{PMI}$ values to 0. To identify clusters in the task space, we embed the relatedness network in a 5-dimensional space using UMAP (McInnes et al., 2018). Next, we cluster sets of contiguous tasks using HDBSCAN (McInnes et al., 2017). These clusters are represented in colors of Fig. 1b. The nodes’ coordinates in this figure are determined by a UMAP embedding of the relatedness network in a 2-dimensional space.
4.5 Task Values
We estimate the value of a task based on salary information from the Stack Overflow developers survey. We focus on the 12,000 US participants that report a salary in 2023. Respondents list the technologies they use, which often correspond to tags in the SO platform. For each SO user, we use the top 300 most similar survey respondents in terms of overlap in tags and technologies to estimate user salary. We use the weighted average salary reported by the most similar respondents, weighting by tag similarity to the SO user. To assign values to tasks, we then calculate the weighted average salary of the users that carry out that task, using the prevalence of the task in users’ task vectors for the years 2018-2023 (the period in which we collect job ads) as weights:
$V_{\theta}$ $=\sum_{u}\frac{X_{u,\theta}^{2018,2023}}{\sum_{u^{\prime}}X_{u^{\prime},\theta}^{2018,2023}}V_{u}$ (3)
$V_{u}$ $=\sum_{\rho\in R_{u}}\frac{M_{u,\rho}}{\sum_{\alpha}M_{u,\alpha}}V_{\rho},$ (4)
where $M_{u,\rho}$ is the number of overlapping tags between SO user $u$ and one of the 300 closest survey respondents, $\rho\in R_{u}$, and $V_{\rho}$ is the salary of survey respondent $\rho$, such that $V_{u}$, the match-weighted average salary of survey respondents matched to user $u$, can be interpreted as a proxy for the user’s salary.
4.6 Job Advertisements
The data on job ads come from the website HackerNews, and were scraped from the aggregator site hnhiring.com. They represent all 50,998 job ads posted to the website between January 2018 and May 2024. These ads are unstructured job descriptions sometimes including salary and expected work activities (tasks). We prompted an LLM (ChatGPT-3.5) to extract salary information and task requirements, which yields about 46,000 jobs listing at least 3 tasks, and 3,949 jobs that contain at least 1 task and salary information. When ads describe more than one position, the LLM provides information on each job separately.
We first relate the task requirements of job ads to the 237 SO tasks. We embed both job ad tasks and SO tags in a 384-dimensional vector space using sBERT (Reimers and Gurevych, 2019). To describe an SO task in this vector space, we calculate the embedding of its label. To increase accuracy, we use somewhat longer labels of up to 30 words generated by ChatGPT-4.0. Finally, we match each task requirement of a job ad to the closest SO tasks based on cosine similarity between the embeddings of the job-ad task and of the SO task. When we cannot find any SO task with a cosine similarity of at least 0.3, we drop the task requirements. This yields a 237-dimensional vector of zeros and ones that describe each job’s requirements in terms of our SO tasks. In Appendix C.3, we demonstrate robustness of our findings to alternative methods and thresholds.
We use these data in two ways to test the validity of our SO tasks. First, we ask whether we can predict the task composition of a job. To do so, we take the SO-task vectors associated with each job ad and stack these vectors into a dataset that contains for each combination of an SO task and an advertised job whether or not the task is required by the job. Next, we mask 40% of observations in this dataset and calculate the density of required tasks in the job around the masked tasks:
$\mathrm{D}_{\theta,j}=\bar{\mathrm{R}}_{\theta,j}\mathrm{T}_{j},$ (5)
where $\bar{}$ denotes row-normalization and $\mathrm{T_{j}}$ and $\mathrm{R}_{\theta,j}$ are the unmasked portions of job $j$’s task vector and the corresponding rows of relatedness matrix $\mathrm{R}$, namely $\mathrm{R}_{\theta,i}=\frac{\mathrm{R}_{\theta,i}}{\sum_{\tau}\mathrm{R}_{\theta,\tau}}$. $\mathrm{D}_{\theta,j}$ denotes the average relatedness of (masked) task $\theta$ to all unmasked tasks in job $j$.
We plot how the estimated probability that a job requires a particular (masked) task changes with that task’s density in the job’s unmasked task requirements. To do so, we bin all masked observations in 10 equally sized bins and calculate relative frequencies with their 95% confidence intervals. These estimated probabilities are plotted on the vertical axis of Fig. 2b, against the bin’s average task density on the horizontal axis.
Our second validation focuses on the set of advertised jobs that also post salary information. It estimates the association between the average value of a job’s SO tasks and the posted wage in the job ad. We report on this in SI, section C.2.
4.7 Analysis of SO users
To study user behavior, we randomly divide the population of users into two equally sized samples. We use the first sample, $S_{1}$, to construct a task relatedness matrix from co-occurrence patterns of tasks in users. The second sample, $S_{2}$, is used to estimate how relatedness affects user activity on SO. To do so, we collect information from all answer posts by these users within rolling two-year intervals. That is, we construct experience vectors $\vec{T}_{u}^{t}$, where $t$ denotes an interval that stretches across the two calendar years preceding year t. For instance, $\vec{T}_{u}^{2018}$ refers to the answers user $u$ provides between January 1, 2016 and December 31, 2017.
4.7.1 Task diversification
Analogously to eq. (5), we use sample $S_{2}$ to calculate for a given user, $u$, and year $t$, the experience density around each task:
$\mathrm{D}^{t}_{u}=\bar{\mathrm{R}}_{S_{1}}\mathrm{T}^{t}_{u},$ (6)
where $\bar{\mathrm{R}}_{S_{1}}$ is the row-normalized relatedness matrix, constructed using users in sample $S_{1}$.
Next, we ask which new tasks users are most likely to enter. To do so, we take all tasks in which user $u$ had no activity in the two calendar years preceding $t$. These tasks represent the set of tasks that user $u$ can enter. We tag all such tasks with a 1 whenever the user posts at least one answer related the the task in year $t$, and with 0 otherwise. Next, we divide all user-task combinations at risk of entry into 10 equally sized bins and estimate the probability that users do enter these tasks in year $t$. Figure 3b shows how these estimated probabilities rise with the user’s experience density around a task. 95% confidence intervals are approximated with a normal distribution, using the following estimate of the standard error of the mean:
$\hat{\sigma}_{b}=\frac{1}{\sqrt{n_{b}}}\sqrt{\hat{\pi}_{b}(1-\hat{\pi}_{b})},$ (7)
where $n_{b}$ the number of observations in bin $b$ and $\hat{\pi}=\frac{d_{b}}{n_{b}}$, where $d_{b}$ the number of entered tasks in bin $b$, the relative frequency with which tasks in bin $b$ are entered.
4.7.2 Voting
To study of a user’s proficiency and learning in a task, we analyze the votes the user’s answers receive. Votes are social feedback from other users indicating appreciation for a post. However, the sample of answers that users post is highly selected. First, before deciding to post an answer, users will look at existing answers to a question to see if they can improve or add to them. To avoid sample selection bias from this, we limit ourselves to the first recorded answers for each question. Second, users will post answers to question for which they perceive themselves to be sufficiently knowledgeable. We address this type of sample selection bias in Appendix D.2, where we use variation in the user base’s activity patterns across the 24-hour daily cycle.
To determine the importance of users’ task experience, we estimate its effect on the upvotes their answers receives. To do so, we estimate the following regression model, with outcomes reported in Fig. 3c:
$y_{a,\theta}=\beta_{x}\log(X_{\theta,u(a)}+1)+\beta_{a}\log(A_{q(a)})+\beta_{v}\log(1+\sum_{i\in Q_{q(a)}}v_{i})+\eta_{a},$ (8)
where $X_{\theta,u(a)}$ is user $u(a)$’s experience in terms of the number of answers provided to questions involving task $\theta$ in the two calendar years preceding the calendar year in which answer $a$ was provided, $A_{q(a)}$ the number of answers provided to question $q(a)$, $Q_{a}$ the set of answers provided to question $q(a)$, $v_{i}$ the number of upvotes that answer $i$ received and $\eta_{a}$ a disturbance term. Furthermore, the dependent variable $y_{a,\theta}$ is either a binary variable that indicates whether answer $a$ is a question’s top answer or $\log(v_{a}+1)$, the logarithm of the number of upvotes answer $a$ receives, augmented by 1 to avoid $\log(0)$ issues. Note that if an answer is associated with multiple tasks, we replicate the observation accordingly.
4.7.3 Task adoption analysis
To study which new tasks users adopt, we estimate the effect of task value and density on new task entries, estimating the following regression model:
$y_{u,\theta}=\beta_{V}\log(V_{\theta})+\beta_{D}D_{u,\theta}+\sigma_{u,\theta},$ (9)
where $y_{u,\theta}$ is a binary variable that encodes whether or not user $u$ adopted task $\theta$, $\log(V_{\theta})$ is the logarithm of the task’s value as defined in eq.( 3), and $D_{u,\theta}$ the user $u$’s density around task $\theta$ as defined in eq. (6) before. We estimate this model once with data from all programming languages and once with data on answers related to Python only. Findings are summarized in Fig. 3d, with detailed results described in Appendix D.3.
Acknowledgements
XF, JW, SD and FN acknowledge financial support from the Austrian Research Agency (FFG), project #873927 (ESSENCSE). JW acknowledges support from the Hungarian National Scientific Fund (OTKA FK 145960). SD acknowledges financial support from the Dutch Research Council (NWO) under the REWIRE project, grant number KICH1.MV02.22.010.
Appendix A Stochastic block models
SBMs represent generative models for networks, that coarsen networks by modeling groupings of nodes and the edge distributions between and within these groupings. These models typically postulate a stochastic data generating process of the network that produces communities and the edges among nodes within and between these communities. Under this assumption, the observed network can be described as sampled from this process.
More precisely, given an $N$-nodes network, $G$, and partition or community structure ${\bf c}$ assigning nodes to different groups, the probability of generating the graph $G$ from its partition ${\bf c}$ is:
$P(G|\bf c,\theta),$ (10)
where $\theta$ is a set of parameters that govern how nodes connect to one another, depending on the communities they belong to. The inference procedure inverses this relation, such that it recovers the probability that a specific generative model is correct, given the observed network $G$. Thus, the probability of observing the configuration given observed network $G$ with model parameters that observe prior probabilities $P(\theta,{\bf c})$ is:
$P({\bf c}|G)=\frac{\sum_{\theta}P(G|\theta,{\bf c})P(\theta,{\bf c})}{P(G)},$ (11)
Under the assumption that the generative model has only one set of particular parameters for network $G$, the equation simplifies to:
$P({\bf c}|G)=\frac{P(G|\theta,{\bf c})P(\theta,{\bf c})}{P(G)}.$ (12)
In the bipartite implementation of this model, priors are chosen such that nodes of the same layer cannot link to one another Gerlach et al. [2018]. Applying this bipartite SBM to our question-tag network yields communities for both questions and tags. In this research, we are only interested in communities of tags (as opposed to questions) of which we discover 247, containing a total of 5,083 tags.
The procedure yields a community label for each tag, even when there is some ambiguity over which community fits the tag best. For instance, some tags are highly ubiquitous (e.g., pandas in Python related questions) and therefore connect to nodes in various communities. This leads to noise in the community composition. To reduce this noise, we try to remove tags that are only weakly connected to the other tags in their community. In principle, SBM models can identify such nodes using MCMC simulations. However, the size of our question-tag renders this approach computationally infeasible. Therefore, we here use a simpler approach.
In particular, we construct a weighted unipartite tag network, where edge weights denote the intensity with which tags co-occur on the same questions. Next, we label nodes in this tag network according to the community structure we discovered above. For each tag, we now calculate the overrepresentation, $O_{tc}$ of (weighted) links to nodes in each of the 247 communities:
$O_{tc}=\frac{w_{tc}/\sum_{c^{\prime}}w_{tc^{\prime}}}{\sum_{t^{\prime}}w_{t^{\prime}c}/\sum_{t^{\prime},c^{\prime}}w_{t^{\prime}c^{\prime}}},$ (13)
where $w_{tc}$ is the sum of edge weights across all links between tag $t$ and community $c$.
$O_{tc}$ measures how important tag $t$ is to community $c$ compared to all other communities. High values of $O_{ct}$ indicate that the tag is a core part of the community, while low $O_{ct}$ values signal that the tag may be misplaced or only weakly related to community $c$. To focus on core tags, for each community, we remove the $20\%$ of tags with the lowest $O_{ct}$ values. Finally, we also remove communities with fewer than 3 tags, for which the tag composition provides too little information on the associated task. In this way, we arrive at a task space with 237 tasks, composed of a total of 4,054 tags.
Appendix B Task space
In this section, we provide a detailed, annotated version of the task space depicted in Fig. 1b. This annotated task space is shown in Fig. B.1.
Each node is labeled with an example of a programming task provided by ChatGPT-4.0. To do so, we prompt the LLM as follows:
I have a list of tags related to software engineering and computer science. The tags are ordered by their importance, listing the most important task first. Using no more than 10 words, could you please suggest a general task related to the following tags ’task_description’ as key words? Please make sure that the task is general and not too specific. Please only return json content and not anything else.
Figure B.1: Fully annotated task space.
Appendix C Analyzing job advertisements
C.1 Example job ad
The analyses in Fig. 2 of the main text, we assign SO tasks to job advertisement. To do so, we first feed the job ad’s text to a large language model (ChatGPT-3.5) to extract strings that describe all task requirements from the listed job description. Next, we use sBert[Reimers and Gurevych, 2019] to embed each of the job ad’s task requirements in an 384 dimensional space. We then do the same for SO task labels we generated in Sec. B. To increase the accuracy of these embeddings we prompt ChatGPT-4.0 to produce somewhat longer labels of up to 30 words. Finally for each task requirement, we pick the closest SO task in terms of the cosine similarity between the two embeddings. We keep only those task requirements for which the matched SO task’s cosine similarity exceeds 0.3.
Below, we illustrate this using a concrete example of a job advertisement. The job advertisement contains the following text:
XXX is a home security and automation company that combines industry-leading technology and truly personal service to protect our customers. We’re looking for a Devops Engineer to work alongside our small engineering team (currently around 6, depending on your definitions) building robust systems and processes. We’d like someone with strong AWS chops and experience with provisioning/config management tools (we currently use Ansible and Terraform). Some other nice points to have would be experience with managing monitoring or metric aggregation systems and PCI compliance. One notable feature to me is the transparency in the organization. The whys and hows of both the past and future are well communicated across most teams. The things we work on are clearly tied to goals, and …
The task requirements the LLM extracts, together with the SO tasks we match to them, are listed in Table C.1. The procedure leads to plausible matches. The first job requirement, AWS (Amazon Web Services, a cloud computing service), is clearly closely related to the matched SO task, which contains several tags related to scheduling tasks, such as job-scheduling or quartz-scheduler and several amazon-specific tags such as aws-application-load-balancer and amazon-athena. Also for the second job requirement, the identified match is plausible. Ansible and Terraform are two platforms for CI/CD (Continuous Integration and Continuous Delivery/Deployment) pipelines, which help automate the building, testing, and deployment of code. In this context, Maven is primarily used in Java and Kubernetes focuses on automating the deployment of containerized applications. For the third and fourth job requirements, the similarity to their matched SO tasks falls below the threshold value of 0.3. Yet, the job requirements and matched SO tasks are still related, albeit somewhat more weakly. In the third job requirement, metric aggregation systems are often used to monitor CPU or network loads, which support scalable real-time data processing. Similary, the PCI (Payment Card Industry compliance mentioned in the fourth job requirement covers security standards required when handling debit or credit card information. This typically involves the implementation of secure user authentication and authorization processes of the matched SO task.
Table C.1: Tasks requirements in job ads and their matching SO tasks.
(1) task requirements (2) matched SO task similarity (cosine)
AWS Deploy scalable cloud-based application using AWS services 0.7987
provisioning/config management tools (Ansible and Terraform) Configure CI/CD pipeline with Kubernetes and Maven 0.4285
managing monitoring or metric aggregation systems Implement scalable real-time data processing pipelines 0.2186
PCI compliance Implement secure authentication and authorization in applications 0.2487
Column (1): required tasks as extracted by the LLM from the job ad. Column (2): the label of the matched SO tasks. Column (3): cosine similarity between the embedding vectors of Columns (1) and (2).
C.2 Salary versus task value
To assess how well SO tasks predict salaries in job ads, we focus on the subset of job advertisements that also provide salary information. For these jobs, we calculate the total number of SO tasks they list, as well as the average value of these tasks. Next, we estimate the following regression model:
$\log(w_{j})=\beta_{v}\bar{v}_{j}+\beta_{r}\bar{r}_{j}+\beta_{n}\log(n_{j})+\tau_{y(j)}+\varepsilon_{j},$ (14)
where $w_{j}$ is the salary or midpoint of the salary range posted for job $j$, $\bar{v}_{j}$ is the average of the logarithm of the task values across all SO tasks job $j$ requires, $\bar{r}_{j}$ the average of the task density around these SO tasks (i.e., the coherence of the tasks the job requires) and $n_{j}$ the number of SO tasks job $j$ requires. $\tau_{y(j)}$ represents year fixed effects to control for general trends in the labor market. Finally, $\varepsilon_{j}$ is an error term.
Outcomes are reported in the table of Table C.2. The estimated effect effect of task value on posted salaries is around 0.9 across all models, suggesting that a $10\%$ increase in task value is associated with a $9\%$ higher salary.
Table C.2: Salary regression in online job ads.
Dependent Variable: $\log(salary)$
Model: (1) (2) (3) (4) (5) (6)
Variables
task value 0.8734∗∗∗ 0.8522∗∗∗ 0.8853∗∗∗ 0.8669∗∗∗
(0.1594) (0.1550) (0.1591) (0.1550)
task coherence 0.0775∗∗ 0.0273 0.0253
(0.0294) (0.0257) (0.0262)
# tasks -0.0059 -0.0067 -0.0066
(0.0034) (0.0032) (0.0033)
Fixed-effects
year Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 4,116 4,116 4,116 4,116 4,116 4,116
R2 0.12595 0.11509 0.12610 0.11501 0.12745 0.12756
Within R2 0.01367 0.00142 0.01384 0.00133 0.01536 0.01548
Robust standard errors in parentheses. p-value: ***: 0.01, **: 0.05, *: 0.1. The dependent variable is the value or midpoint of the salary range posted in a job ad, task value is the average of the logarithm of the task values across all SO tasks the job requires, task coherence is the average of the task density round all SO tasks the job requires, # tasks is the logarithm of the number of SO tasks the job requires.
C.3 Alternative job task matches and thresholds
We test three alternative methods to match task requirements in job ads to SO tasks. To do so, let $v_{t}^{j}$ denote the embedding vector of task requirement $t$ of job $j$. Our first method, used in the main text, calculates the cosine similarity between $v_{t}^{j}$ and the embedding vector of each of the SO task’s (long) labels, and then selects the SO task that yields the highest cosine similarity. Our second method uses the similarity of $v_{t}^{j}$ to the embedding of the most prominent tag (in terms of its usage on SO) in each SO task. Our third method calculates the similarity of $v_{t}^{j}$ to the average embeddings of the tags defining the SO task. Finally, our fourth method uses the similarity of $v_{t}^{j}$ to the embedding of the closest tag in each SO task. Next, we repeat the analysis depicted in Fig. 2 with each of these methods. Figure C.1 presents the results.
Figure C.1: Alternative methods to match task requirements in job ads to SO tasks. a,c,e,g. Prediction of task requirements. $20\%$ of task vector elements are masked and grouped into 10 equally sized bins based on the average relatedness of the masked task to the (unmasked) SO task requirements of a job. Plot displays the estimated probability, with its $95\%$ confidence interval, that tasks in a bin are required in the job. b,d,f,h. Predictions of wage offers. Jobs are grouped into equally sized bins based on the average value of the tasks they require. Vertical axis shows the estimated mean salary posted in the job ads, with their $95\%$ confidence intervals. a,b. Task requirements are matched to SO tasks using the embedding vector of the SO task’s label, namely the results shown in Fig 2. c,d. Task requirements are matched to SO tasks using the embedding vector of the SO task’s main (most frequent) tag. e,f. Task requirements are matched to SO tasks using the average embedding vector across all of the SO task’s tags. g,h. Task requirements are matched to SO tasks using the closest embedding vector across all tags of SO tasks.
In the main text, we only keep task requirements that can be matched with SO tasks at a cosine similarity of $0.3$ or above. To explore the significance of this threshold, we repeat our analyses with two another thresholds: $0.2$ and $0.4$. The results are shown in Fig. C.2. Varying the threshold around the $0.3$ value of the main text does not meaningfully alter the outcomes of this analysis: density around masked tasks is strongly predictive of jobs demanding these tasks and salaries posted in job ads rise with the average values of the SO tasks we identify for them.
Figure C.2: Results of task and salary prediction at varying matching thresholds. a,c,e. Probability that a masked SO task is listed among a job’s task requirements. Masked tasks represent $40\%$ of task vector elements and grouped into 10 equally sized bins based on the average relatedness of (density around) the masked task to the (unmasked) task requirements of a job. Plot displays estimated probabilities in each bin. b,d,f Predictions of salaries. Jobs are grouped into equally sized bins based on the average value of their required matched SO tasks. Vertical axis shows the estimated means in each bin of the advertised wage offers. a,b. Task requirements are matched to SO tasks with a minimum cosine similarity of $0.3$, namely the results shown in Fig. 2. c,d. Task requirements are matched to SO tasks with a minimum cosine similarity of $0.2$. e,f. Task requirements are matched to SO tasks with a minimum cosine similarity of $0.4$. Vertical bars represent 95% confidence intervals.
Appendix D Analysis on SO users
D.1 Alternative specifications for voting regression
In Fig. 3c of the main text, we study how the popularity of an answer relates to the task experience of the user that provided it. The figure shows that answers by users with more task experiences tend to receive more upvotes and are more likely to be the top answer to a question. Table D.1 provides the full set of regression results related to this analysis.
Table D.1: Voting regressions.
Dependent Variables: top answer (binary) # votes (log)
Model: (1) (2) (3) (4)
Variables
task experience 0.0097∗∗∗ 0.0104∗∗∗ 0.0130∗∗∗ 0.0137∗∗∗
(0.0008) (0.0002) (0.0016) (0.0003)
# answers -0.4886∗∗∗ -0.4915∗∗∗ -0.7939∗∗∗ -0.7982∗∗∗
(0.0036) (0.0019) (0.0199) (0.0037)
# votes all answers -$0∗∗$0402∗∗∗ 0.8604∗∗∗ 0.8596∗∗∗
(0.0023) (0.0006) (0.0037) (0.0011)
Fixed effects
year Yes Yes
year-task Yes Yes
Fit statistics
Observations 7,984,391 7,984,391 7,984,391 7,984,391
R2 0.28841 0.28936 0.81479 0.81518
Within R2 0.27190 0.26689 0.80879 0.80359
p-value: ***: 0.01, **: 0.05, *: 0.1
Regression model of eq. (8). The dependent variable is either a binary variable that encodes whether or not an answer has received the most upvotes among all answers provided to a question, or the logarithm of the number of upvotes the answer received. task experience: logarithm of task experience, the total number of answers the users provided to questions related to the focal task in the past two calendar years. # answers: logarithm of the total number of answers provided to the focal answer’s question. # votes all answers: logarithm of the total number of votes received by these answers. Regression models contain either year or task-year fixed effects.
Because for questions that receive only a single answer, this answer is automatically the top answer, we repeat the analysis in a sample that excludes any question that receives only one answer. Table D.2 shows that this leads to larger observed effects of task experience.
Table D.2: Voting regressions when questions receive at least two answers.
Dependent Variables: top answer (binary) # votes (log)
Model: (1) (2) (3) (4)
Variables
task experience 0.0260∗∗∗ 0.0279∗∗∗ 0.0313∗∗∗ 0.0336∗∗∗
(0.0009) (0.0004) (0.0026) (0.0007)
# answers -0.2437∗∗∗ -0.2491∗∗∗ -0.6309∗∗∗ -0.6371∗∗∗
(0.0043) (0.0021) (0.0216) (0.0036)
# votes all answers -0.0814∗∗∗ -0.0816∗∗∗ 0.7513∗∗∗ 0.7503∗∗∗
(0.0065) (0.0011) (0.0046) (0.0014)
Fixed-effects
year Yes Yes
task-year Yes Yes
Fit statistics
Observations 3,170,868 3,170,868 3,170,868 3,170,868
R2 0.09076 0.09334 0.67850 0.67968
Within R2 0.08081 0.08076 0.66944 0.66110
p-value: ***: 0.01, **: 0.05, *: 0.1
Idem Table D.1, but excluding answers to questions that receive only a single answer.
D.2 Causal effects: Instrumental variable estimation
In the main text, we raise the issue of sample selection bias in the regression analysis of the number of upvotes an answer receives on SO. There are two concerns. First, users can observe all existing answers to a question and, arguably, users will take these answers into consideration before deciding whether or not to provide an answer themselves. As a consequence, the answers we observe on SO are no random sample of user-answer combinations. For instance, we expect that users who believe they have little to add to the existing answers to a given question will tend to refrain from posting their own answers. This bias is addressed by only focusing on the first answer (in a temporal sense) that each question received.
Second, even the first answer to a question will not be randomly sampled from a population of users and the answers they would have provided had they chosen to do so. Instead, answers will be provided by users who believe they have useful insights to the problem at hand. To address the sample selection bias this causes, we will exploit the fact that, due to timezone differences, most users are active on the SO platform at somewhat predictable times. Therefore, we expect that who will provide the first answer to a question to some extent depends on the time of day the question was posted. This introduces some exogenous variation in the likelihood that a given user provides an answer to a given question. Moreover, as far as task experience is not fully equally distributed across timezones, this offers exogenous variation in the task experience associated with the first answer to a question.
We proceed as follows. First, we label each minute of a day, henceforth day-minute, from 0 to 1399 in a manner that is independent of timezones. Next, for each user, we calculate the average minute of the day they post answers on SO. This offers a coarse indication of when a user is active on the platform. Finally, we calculate for each task and each day-minute the amount of task experience that is expected to be available on SO. To do so, we sum the task experience of all users, weighted by the gap between the day-minute of a question and the average day-minute of each user. We call this variable, $M_{\theta,m}$, the minute-task count of task $\theta$ at day-minute $m$. Next, we use this variable as an instrument for the actual task experience of whichever user provides the first answer to the question at hand.
Note that there are some potential concerns about this instrument. First, if there are geographic components to task experience of users and if these geographic components directly influence the quality of an answer, the instrument may be invalid. For instance, the timezone that includes Silicon Valley may have a disproportionate number of highly skilled programmers, such that any questions posed when most people in this timezone are active may attract high quality first answers. To correct for this, we always include day-minute fixed effects and in some specifications also add task-year-combination fixed effects.
Second, the instrument assumes that questions are answered within day from the time they are posted on SO. In fact, we would expect that our instrument does not work for questions whose first answer does not arrive within 24 hours. We exploit this in placebo tests that focus on questions whose first answer is provided at least 24 hours after the posting of the question itself.
Table D.3 shows results for an adapted version of the regression model of eq. (8) in the main text, but now using Two-Stage Least Squares (2SLS) estimation. In particular, we estimate the following equation:
$y_{a,\theta}=\beta_{x}\log(X_{\theta,u(a)}+1)+\beta_{a}\log(A_{q(a)})+\beta_{v}\log(1+\sum_{i\in Q_{q(a)}}v_{i})+\mu_{q(a)}+\eta_{a},$ (15)
where $u(a)$ is the user who posts answer $a$, $X_{\theta,u(a)}$ user $u(a)$’s experience in terms of the number of answers provided to questions involving task $\theta$ in the two calendar years preceding the calendar year in which answer $a$ was provided, $A_{q(a)}$ the number of answers that will eventually be provided to question $q(a)$, $Q_{a}$ the set of these answers, $v_{i}$ the number of upvotes that answer $i$ received and $\eta_{a}$ a disturbance term. Compared to eq. (8) in the main text, we add day-minute fixed effects, $\mu_{(q_{a}})$, that control for any timezone specific confounders.
As before, the dependent variable $y_{a,\theta}$ is either a binary variable that indicates whether answer $a$ is the top answer to its question or $\log(v_{a}+1)$, the logarithm of the number of upvotes answer $a$ receives, augmented by 1 to avoid $\log(0)$ issues. If an answer is associated with multiple tasks, we replicate the observation accordingly.
The results is presented in Table D.3. The effect of task experience on answer popularity is significant and positive across all model specifications. Moreover, the estimated causal effects arrived at by instrumental variables estimation are substantially larger than the OLS estimates reported in the main text. This is in line with a downward sample-selection bias in the OLS results.
Table D.3: Instrumental variable regression.
Dependent Variables: top answer (binary) # votes (log)
Model: (1) (2) (3) (4)
Variables
task experience 0.0162∗∗∗ 0.0601∗∗∗ 0.0306∗∗∗ 0.0919∗∗∗
(0.0005) (0.0179) (0.0006) (0.0209)
# answers -0.4847∗∗∗ -0.4598∗∗∗ -0.7835∗∗∗ -0.7484∗∗∗
(0.0006) (0.0114) (0.0008) (0.0133)
# votes all answers -0.0418∗∗∗ -0.0536∗∗∗ 0.8555∗∗∗ 0.8385∗∗∗
(0.0002) (0.0048) (0.0003) (0.0056)
Fixed-effects
year Yes Yes
QA-abs-minute Yes Yes Yes Yes
task-year Yes Yes
Fit statistics
Observations 7,984,391 7,984,391 7,984,391 7,984,391
R2 0.28734 0.22128 0.81305 0.78261
Within R2 0.27051 0.19626 0.80680 0.76881
Clustered (group_task_qminute) standard-errors in parentheses
Signif. Codes: ***: 0.01, **: 0.05, *: 0.1
Results of eq. (15) estimated with 2SLS. The dependent variable is either a binary variable that encodes whether or not an answer has received the most upvotes among all answers provided to a question, or the logarithm of the number of upvotes the answer received. task experience: logarithm of task experience, the total number of answers the user provided to questions related to the focal task in the past two calendar years. # answers: logarithm of the total number of answers provided to the focal answer’s question. # votes all answers: logarithm of the total number of votes received by these answers. Regression models contain day-minute fixed effects (QA-abs-minute) for the timing of the question and year or task-year fixed effects. To overcome sample-selection biases, we instrument task experience with $M_{\theta,m}$, the estimated amount of available task experience at the moment the question was posted, calculated as the weighted sum of task experience across all users, where weights reflect the temporal distance between the day-minute of a user and of the question’s posting.
Next, we test this strategy using our placebo tests. To do so, we split the sample of questions into questions that receive a first answer within the first 24 hours after they were posted and those that receive their first answer later than that. Results are shown in Table D.4. For answers provided within 24 hours, task experience has a strong and statistically significant positive effect on answer popularity, regardless of whether we look at the probability of being the top answer or the number of upvotes an answer receives. In contrast, for the placebo test, comprising of answers provided more than 24 hours after the question was posted, the effect of task experience becomes insignificant in three out of four models. In the model without task fixed effects, the effect on the number of votes an answer receives turns significantly negative. This suggests that the task fixed effects control for some relevant confounders. Therefore, our preferred estimates are derived from models that include these fixed effects. These models suggest a large and positive effect of task experience on the popularity of answers, corroborating our analysis in the main text.
Table D.4: Instrumental variable regression: placebo tests.
Dependent Variables: top answer (binary, $\leq\text{24 hours}$) # votes (log, $\leq\text{24 hours}$) top answer (binary, $\geq\text{24 hours}$) # votes (log, $\geq\text{24 hours}$)
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
task experience 0.0183∗∗∗ 0.0620∗∗∗ 0.0314∗∗∗ 0.0947∗∗∗ 0.0009 0.0561 -0.0172∗∗∗ -0.0030
(0.0005) (0.0174) (0.0006) (0.0203) (0.0038) (0.2365) (0.0048) (0.2827)
# answers -$0∗∗$4528∗∗∗ -0.7750∗∗∗ -0.7347∗∗∗ -0.5116∗∗∗ -0.4920∗∗∗ -0.8835∗∗∗ -0.8773∗∗∗
(0.0007) (0.0123) (0.0008) (0.0143) (0.0023) (0.0848) (0.0032) (0.1014)
# votes all answers -0.0448∗∗∗ -0.0573∗∗∗ 0.8471∗∗∗ 0.8284∗∗∗ -0.0231∗∗∗ -0.0275 0.9150∗∗∗ 0.9140∗∗∗
(0.0002) (0.0050) (0.0004) (0.0058) (0.0004) (0.0190) (0.0007) (0.0227)
Fixed-effects
year Yes Yes Yes Yes
QA-abs-minute Yes Yes Yes Yes Yes Yes Yes Yes
task-year Yes Yes Yes Yes
Fit statistics
Observations 6,915,424 6,915,424 6,915,424 6,915,424 1,068,555 1,068,555 1,068,555 1,068,555
R2 0.28071 0.21444 0.80404 0.77090 0.32491 0.24001 0.86829 0.87025
Within R2 0.26317 0.18891 0.79722 0.75538 0.31531 0.22360 0.86502 0.86413
Clustered (group_task_qminute) standard-errors in parentheses
p-value: ***: 0.01, **: 0.05, *: 0.1
Idem Table D.3, but splitting the sample into questions that were answered within 24 hours and those that weren’t. The results for the latter sample, depicted in the in columns (5)-(8), represent placebo tests for the instrumental variables estimation.
D.3 Regression of entry on task value
To study how task values effect users’ adoption on new tasks, we run the regression described in eq. (9) of the main text. We compare the association of two variables, task value and task density, with the likelihood that users adopt a certain task. We repeat this analysis once using all answers to construct experience and task entry variables, and once using only answers to Python related questions. The results are shown in Table D.5.
Table D.5: Task entry regressions
Dependent Variable: entry (binary, all programming languages) entry (binary, Python)
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
task value -0.0078∗∗∗ -0.0079∗∗∗ -0.0068∗∗∗ -0.0069∗∗∗ 0.0342∗∗∗ 0.0348∗∗∗ 0.0140∗∗∗ 0.0146∗∗∗
(0.0008) (0.0008) (0.0006) (0.0006) (0.0011) (0.0011) (0.0009) (0.0011)
density 0.0194∗∗∗ 0.0173∗∗∗ 0.0181∗∗∗ 0.0176∗∗∗
(0.0002) (0.0002) (0.0004) (0.0006)
Fixed-effects
user Yes Yes Yes Yes
Fit statistics
Observations 24,367,254 24,367,254 24,367,254 24,367,254 2,229,882 2,229,882 2,229,882 2,229,882
R2 $1.8\times 10^{-5}$ 0.01688 0.01859 0.02743 0.00047 0.01627 0.02210 0.03082
Within R2 $1.86\times 10^{-5}$ 0.01076 0.00049 0.01528
Detailed outcomes of the regression model described in eq. (9) of the main text. Observations are user-task combinations for which the user did not answer any questions related to the task in the previous two years. The dependent variable is a binary value of whether or not the user enters the task (i.e., answers questions related to this task) in the current period. task value: logarithm of the task’s imputed value, task density: the user’s prior experience in other tasks, weighted by the relatedness of these tasks to the current task. For ease of interpretation, we subtract the variable’s mean and then divide by its standard deviation. Models in columns (2), (4), (6) and (8) control for user fixed effect. Columns (1)-(4) report analyses performed using all answer posts, columns (5)-(8) use only Python related answer posts. Clustered (user) standard-errors in parentheses. p-values: ***: 0.01, **: 0.05, *: 0.1.
Appendix E Rescaling by Github
Figure 4 shows Python’s progressive rise as the dominant language for an increasing number of programming tasks. However, even though SO may provide a useful dataset to define such programming tasks, it is not guaranteed to be representative of programmers across the world. To assess this problem, we compare counts of programmers on SO to comparable ones on GitHub. GitHub is the world’s largest software development collaboration platform and therefore offers a good representation of the global software development sector. In particular, we collect information on the (self-reported) countries of residence on GitHub and (observed) language use between 2020 and 2022 [GitHub, 2025]. Figure E.1a compares SO to GitHub user counts at the level of countries, languages and country-language combinations. This shows that user numbers are highly correlated across the two platforms, with correlations ranging from 0.75 to 0.93.
Figure E.1: a. The GitHub user counts against the SO user counts in the perspectives of country, programming languages and country-language in the year 2022. The values are in log scale and the $R^{2}$ values of the regression fit are provided inside each panel. b. Number of tasks in which a programming language is the top language in terms of SO users reweighted by user counts from GitHub, from 2020 to 2022. The graph shows the largest 6 languages measured by cumulative SO posts between August 2008 and June 2023.
Nevertheless, some languages are better represented on SO than others. This may affect our analysis of how dominant a language is for a given task. In Fig. E.1b we attempt to correct for this, by reweighting the user counts by language using their observed shares on GitHub. This reweighting ensures that the distribution of users across language on SO mimics the one on GitHub. Because we only have GitHub information for the years 2020, 2021 and 2022, the analysis is limited to those years. Also with reweighted user counts, Python leads in most tasks, albeit with round 50, in somewhat fewer tasks than what we report in the main text. The ordering of the remaining languages is mostly unchanged as well, except for C#, which was somewhat more prominent before rescaling user counts. Moreover, also the most prominent temporal patterns over the 3-year time period are mostly left unchanged. Figure E.1b retains the rise of Python and the drops of Java and PHP, as well as the relative stability in these years of JavaScript, validating the results shown in the main text.
Appendix F Swift vs Objective-C
Figure F.1: Number of tasks in which Objective-C has more users on SO than Swift and vice versa. The dashed vertical line marks the year 2014, the year that Swift was released.
Software written for Apple devices was mostly written in Objective-C from the 1980s until the mid 2010s. Apple designed Swift to replace Objective-C and launched the language in 2014. Within the next few years, most users switched to this new language to be able to continue developing software for Apple products  [JetBrains, 2023]. Unlike typical programming language dynamics, which evolve organically from developer communities, Swift’s introduction was a deliberate, top-down intervention by Apple.
Figure F.1 describes this shift in terms of the tasks where Swift was more popular than Objective-C and vice versa. Before 2015, most tasks were handled preferentially in Objective-C. However, since the release of Swift in 2014 this rapidly changes. Swift surpasses Objective-C in terms of the number of different tasks it handles around 2015 and all but eclipses Objective-C by 2016.

### v2

@@
-The building blocks of software work explain coding careers and language popularity
+Using digital traces to analyze software work: skills, careers and programming languages
 Abstract
-Recent waves of technological transformation have fueled debates about the changing nature of work.
-Yet to understand the future of work, we need to know more about what people actually do in their jobs, going beyond educational credentials or job descriptions.
-Here we analyze work in the global software industry using tens of millions of Question and Answer posts on Stack Overflow to create a fine-grained taxonomy of software tasks, the elementary building blocks of software development work.
-These tasks predict salaries and job requirements in real-world job ads.
-We also observe how individuals learn within tasks and diversify into new tasks.
-Tasks that people acquire tend to be related to their old ones, but of lower value, suggesting that they are easier.
-An exception is users of Python, an increasingly popular programming language known for its versatility.
-Python users enter tasks that tend to be higher-value, providing an explanation for the language’s growing popularity based on the tasks Python enables its users to perform.
-In general, these insights demonstrate the value of task taxonomies extracted at scale from large datasets: they offer high resolution and near real-time descriptions of changing labor markets.
-In the case of software tasks, they map such changes for jobs at the forefront of a digitizing global economy.
-Classification: Social Sciences/Economics
-Keywords: human capital; tasks; software; economic complexity; networks
+Recent waves of technological transformation are reshaping work in uncertain and hard-to-predict ways.
+However, jobs at the forefront of the digitizing economy offer an early glimpse of these changes and leave rich activity traces.
+We exploit such traces in tens of millions of Question and Answer posts on Stack Overflow for the creation of a fine-grained taxonomy of software skills to analyze human capital in the global software industry.
+Constructing a software skill space that maps relations among these skills reveals that real-world software jobs demand highly coherent skill sets and that programmers learn through a process of related diversification.
+The latter process often leads to the acquisition of lower-value skills.
+However, when programmers use Python they preferentially target higher-value skills, offering a potential explanation for Python’s successful rise as a dominant general purpose language.
+Keywords: human capital; skills; software; economic complexity; networks
 1 Introduction
-Knowledge of the tasks that workers perform is central to understanding wages, productivity, and overall labor market dynamics (Acemoglu and Autor, 2011).
-Recent work has transformed our ability to do so by using complex network analysis — moving away from descriptions that rely on years of schooling, credentials, and work experience to the granular building blocks of human capital: skills and tasks (Anderson, 2017; Alabdulkareem et al., 2018; Hosseinioun et al., 2025).
-These studies focus on jobs, describing them as networks of skills or tasks using data on skill requirements and task descriptions extracted from large-scale surveys (one, 2025), expert assessments (Commission, 2017), job ads (Börner et al., 2018) or online freelance platforms(Anderson, 2017; Stephany and Teutloff, 2024).
-Yet, job descriptions tell us something about the nature of vacancies, not of the people who fulfill them.
-This limits our ability to understand how workers specialize and learn new things, processes critical to a successful career.
-Focusing on the software sector, we propose a new approach that uses data from Stack Overflow (SO), a very large, online question-answer database.
-The core insight we leverage is that questions on SO describe coding problems that programmers encounter in their work.
-As such, they can be understood as instantiations of software-development job-tasks.
-By generalizing these descriptions of specific problems encountered at work, we can construct a taxonomy of canonical software tasks that is highly granular, flexible and based in data.
-Moreover, we can use this taxonomy to describe the human capital endowments of people, not the human capital requirements of jobs.
-That is, we ask what programmers do at work, not what they are supposed to do according to their job descriptions.
-This allows us to analyze how programmers learn new tasks and what the value of these tasks is.
-Moreover, programmers can choose from a wide menu of programming languages and we show that programming languages differ in the tasks they support.
-However, the mapping of tasks to programming languages changes over time, as illustrated by the rapid rise of Python as a general purpose language.
-Our task network shows that this rise exploits a remarkable strength of Python: it allows programmers to more quickly transition to the most valuable programming tasks.
-As wages and careers have come to depend on how workers specialize (Hosseinioun et al., 2025), interact (Deming, 2017), complement colleagues (Neffke, 2019), and move across places and jobs (Frank et al., 2024), fine-grained taxonomies of tasks are increasingly valuable.
-Such taxonomies clarify how workers transition between occupations (del Rio-Chanona et al., 2021), how skills can substitute for or complement each other (Anderson et al., 2012; Neffke, 2019), and how occupations evolve over time (Nedelkoska et al., 2021).
-In this paper, we focus on jobs in the global software development sector.
-Software production is a fast-changing and critical domain of the digital economy (Brynjolfsson and McAfee, 2014; Wachs et al., 2022; Juhász et al., 2024).
-Of the seven largest companies by market capitalization worldwide, three (Microsoft, Alphabet, and Meta) focus almost exclusively on software, while three others (Nvidia, Apple, and Amazon) derive substantial revenue from it.
-Demand for programming skills continues to surge: one in 11 U.S.
-occupations now requires coding (one, 2025), and software-related roles such as data scientist, information security analyst, and computer research scientist rank among the fastest-growing occupations (U.S.
+Knowledge of detailed skills and expertise used at work is central to understanding wages, productivity, and overall labor market dynamics (Acemoglu and Autor, 2011).
+Recent work has transformed our ability to do so by using complex network analysis.
+This has allowed moving away from descriptions that rely on years of schooling, credentials, and work experience to studies of the relations among concrete skills and the tasks they enable workers to perform (Anderson, 2017; Alabdulkareem et al., 2018; Hosseinioun et al., 2025).
+However, most of these studies focus on the content of jobs, not the human capital of who performs them.
+This limits our understanding of how workers learn new skills, which is crucial for helping individuals, firms and educators navigate transforming labor markets in times of rapid technological change.
+To address this, we focus on a specific type of work that has become increasingly important, is changing fast and can be studied in great detail: software development.
+Constructing a fine-grained taxonomy of programming skills, we create a software development skill space that maps relations among these skills.
+We use this space to analyze which skills are combined in job ads, how programmers acquire new skills and what role programming languages play therein.
+Software jobs are particularly interesting, because they represent a fast-changing and critical domain of the digital economy (Brynjolfsson and McAfee, 2014; Wachs et al., 2022; Juhász et al., 2026).
+Software companies dominate the global ranks of the largest firms by market capitalization.
+[^1] Of the seven largest companies worldwide, three (Microsoft, Alphabet, and Meta) focus almost exclusively on software, while two others (Apple and Amazon) derive substantial revenue from it.
+On the labor market, demand for programming skills has risen sharply over the past two decades: one in 11 U.S.
+occupations now requires coding skills (O*NET, 2025), and software-related roles such as data scientist, information security analyst, and computer research scientist rank among the fastest-growing occupations in the US (U.S.
 Bureau of Labor Statistics, 2025).
-As new hardware, programming languages, and specialized libraries reshape computing, fresh roles (e.g., web development, DevOps, data science, artificial intelligence) keep emerging (Rock, 2019; Papoutsoglou et al., 2019; Hemon et al., 2020).
-This changing landscape has sparked intense global competition for talent, prompting many major economies to designate software development as a strategic priority (Draghi, 2024; Kazim et al., 2021; Singh, 2024).
-However, we currently lack a detailed understanding of the fast-changing human capital that is employed in this sector.
-One reason is that databases linking occupations to tasks (e.g., O*NET (one, 2025)) are often curated by human experts, making them relatively coarse and slow to update.
-In O*NET, for example, software-related tasks often focus on broad domains of application (e.g., “geoinformatics” or “biostatistics”), team dynamics (e.g., managing or collaborating with software engineers), or acquiring software products.
-Alternative data sources such as job advertisements or resumes are also challenging because job ads describe job requirements of ideal-type workers, not human capital endowments of actual workers, resumes can be aspirational, and both often focus on limited geographies.
-Likewise, platforms for online freelancing provide insights into how complex skill interdependencies affect wage premiums (Anderson, 2017; Stephany and Teutloff, 2024), but the extent to which these findings generalize to the broader labor market remains uncertain.
-We address these gaps by analyzing millions of posts from SO—the largest question-and-answer forum for programmers (Anderson et al., 2012).
+At the macro-economic level, this has sparked intense global competition for talent, prompting major economies to designate software development as a strategic priority (Draghi, 2024; Kazim et al., 2021; Singh, 2024).
+However, we currently lack a detailed understanding of the changing human capital that is employed in this sector.
+As new hardware, programming languages, and specialized libraries reshape computing, fresh roles (e.g., web development, DevOps, data science, artificial intelligence) keep emerging (Rock, 2019; Papoutsoglou et al., 2019; Hemon et al., 2020) and existing roles transform.
+Existing databases linking occupations to skills (e.g., O*NET, 2025)) are often curated by human experts, making them relatively coarse and slow to update.
+[^2] For example, the only software-related skill in O*NET is undifferentiated “programming” and the listed software-related work activities often focus on broad domains of application (e.g., “geoinformatics” or “biostatistics”), team dynamics (e.g., managing or collaborating with software engineers), or acquiring software products.
+More dynamic data sources are challenging in their own right.
+For instance, job ads — frequently used to study job contents — describe job requirements of ideal-type workers, not human capital endowments of actual workers.
+Likewise, although rich data from platforms that mediate online freelancing work provide insights into how complex skill interdependencies affect wage premiums (Anderson, 2017; Stephany and Teutloff, 2024), the extent to which these findings generalize to traditional labor markets remains uncertain.
+We address these gaps by analyzing millions of posts from Stack Overflow (SO)—the largest question-and-answer forum for programmers (Anderson et al., 2012).
 Each post describes a specific technical challenge and includes tags that indicate the tools, languages, or concepts involved (Barua et al., 2014).
-These challenges reflect real-world coding tasks and users use their posting histories in part to signal their skills to prospective employers (Xu et al., 2020).
-Interpreting tags as a structured, albeit partial, description of posted questions, we aggregate questions to create a software-task taxonomy by identifying clusters of frequently co-occurring tags.
-Using this taxonomy, we examine how software tasks evolve in the labor market.
-We find that task value estimates, derived from survey data, significantly predict salaries in real-world job postings.
-Notably, wages vary greatly among these tasks, with higher-paying ones concentrated in more specialized or in-demand fields such as AI and machine learning.
-We also introduce a notion of task relatedness based on co-occurrence patterns in users’ posting histories.
-This measure of relatedness derived from interactions with the SO platform predicts similar co-occurrence trends in job ads, suggesting that our taxonomy captures a meaningful structure of the software development landscape.
-At the individual user level, we observe two learning patterns.
-First, developers improve in the tasks they gain experience in, as measured by community feedback on their posts.
-Second, developers enter new tasks over their SO careers, and task relatedness strongly influences these entries: developers are fifteen times more likely to move into tasks related to ones they already know.
-At the same time, we find that developers usually add tasks with lower market value, which are presumably easier to master.
-Finally we demonstrate that our fine-grained software task taxonomy sheds light on the evolution of software development as a whole.
-We observe that programming languages differ in the number of tasks for which they are used, following a nested pattern of tasks in programming languages reminiscent of similar patterns observed in ecology (Bascompte et al., 2003), economic development (Hidalgo and Hausmann, 2009; Tacchella et al., 2012; Mariani et al., 2019) and skill hierarchies (Hosseinioun et al., 2025).
-Accordingly, instead of specialization, where different languages are used to perform different types of tasks, we observe that tasks have an implicit hierarchy such that the most ubiquitous tasks can be performed in all languages, whereas rare tasks are only performed by the most versatile ones.
-A prime example of the latter is the Python programming language, a popular language that has seen significant growth in recent years (Economist, 2018).
-We show that this growth has been accompanied by changes in the way Python is used.
-First, Python has become increasingly versatile.
-From 2015, the number of tasks in which Python dominates as the preferred language for most programmers has rapidly grown, rendering Python into one of today’s premier general purpose languages.
-Second, Python enables qualitatively different paths for users to learn new tasks.
-Specifically, unlike users of most other languages, Python users often branch into more, not less valuable tasks.
-This observation could help explain Python’s success, despite its not being the oldest or fastest language: its ecosystem may make it easier for developers to tackle higher-value tasks.
-2 Results
-2.1 Constructing a space of software tasks
-Software development tasks can be viewed as recurring categories of programming problems developers encounter.
-SO, with its vast repository of user-generated questions, offers a unique window into these real-world challenges.
-To organize this repository and facilitate retrieval, each question is labeled with one or more of 66,000 curated tags that indicate the specific tools, frameworks, or concepts involved.
-We use these tag–question relationships to group related issues into “canonical” software tasks, applying a bipartite stochastic block model (SBM) (Peixoto, 2017, 2014; Gerlach et al., 2018) to detect communities in the graph of questions and tags (Fig.
-1a).
-We ignore tags that refer to programming languages and use them later on to study how tasks are distributed across languages.
-To generate concise names for each task community, we prompt ChatGPT-4.0 for labels that capture the essence of the tags comprising each group (see Appendix B).
-This process yields a set of 237 software tasks, such as Develop NLP models and tools for text analysis, Establish secure server access and control, and Deploy and secure a scalable web app.
-Figure 1: Mapping software tasks.
-a.
-Stylized depiction of the bipartite question-tag network.
-SBM groups tags into communities (tasks) that connect to similar sets of questions.
-ChatGPT-4.0 finds a common label that summarizes each community’s tag information.
-b.
-Task space.
-Pointwise mutual information ($\mathrm{PMI}$) expresses how surprisingly often two tasks are performed by the same users.
-UMAP embeds the resulting co-occurrence network in a 2-dimensional plane (the task space).
-c.
-Close-up on Develop AI models task, depicting the original network structure among the 10 most closely related tasks.
-d.
-Task values.
-Nodes are colored according to their task value, estimated from salary information in the SO 2023 developers survey.
-Darker shades indicate more valuable tasks.
-e.
-Table of the five most and least valuable tasks.
-For details, see Methods.
-Because certain tasks complement each other, they often co-occur within the same job roles (Anderson, 2017; Alabdulkareem et al., 2018; Neffke, 2019; Stephany and Teutloff, 2024).
-To investigate this task bundling, we identify which tasks are frequently performed by the same SO users.
-This yields a network that links tasks based on their co-occurrence in user activity.
-To reduce dimensionality and facilitate visualization, we then embed this network in a two-dimensional plane.
-The resulting task space contains several clusters of related tasks, each shown in a different color in Fig.
-1b (see Methods).
-Many of these clusters group tasks with comparable labor market returns.
-This is illustrated in Fig.
-1d, where, we plot task values inferred from self-reported wages in a large-scale survey run among SO users (see Methods).
-On the right side of the task space, for example, three clusters revolve around website development, typically linked to lower wages.
-By contrast, tasks in the top center—focusing on advanced iOS and Android development—offer significantly higher earnings.
-Additional high-value clusters include those for AI and machine learning (AI/ML), Advanced Programming Concepts, Cloud Computing, and DevOps.
-The table of Fig.
-1e lists the five highest-paying tasks, three of which are related to iOS app development, and five lowest-paying tasks, all associated with more basic web development.
-Notably, clusters that are conceptually linked—such as AI/ML and Statistics and Data Analysis, and iOS and Android development—tend to lie near each other in the task space.
-To investigate the network structure of the task space, we can examine individual tasks in greater detail.
-For example, focusing on the ten tasks most closely related to developing AI models, Fig.
-1c reveals strong connections to other statistical tasks, as well as tasks involving image and natural language processing, data visualization, and packaging models into standalone applications.
-2.2 Validation using job ads
-How well do these software tasks capture real-world development work beyond SO?
-To find out, we examine approximately 50,000 software job ads from HackerNews [^1] https://www.hackernews.com/ (see Fig.
-2a and Methods).
-Using an LLM, we parse each job ad’s text to extract salaries and the list of tasks required of successful candidates.
-Next, we convert each job’s extracted tasks into a 237-dimensional binary vector, where each dimension corresponds to one of the SO tasks identified previously.
-We accomplish this by embedding both the job’s task descriptions and the SO task labels into a 384-dimensional space using sBERT(Reimers and Gurevych, 2019).
-We then match each extracted task to its closest SO task by comparing cosine similarities.
-In this way, each job ad becomes a vector of job requirements, expressed in terms of SO tasks.
-Finally, we combine this vector with task-value information and measures of task relatedness.
-Figure 2a (see Methods) illustrates this process.
-Figure 2: Job ads.
-a.
-Schematic representation of the workflow to extract salary and task requirements from online job ads by prompting ChatGPT.
-Task requirements are converted to the 237-dimensional SO task vectors of Fig.
-1 based on cosine similarities between text embeddings of task requirements and SO tasks.
-b.
-Prediction of task requirements.
-40% of task vector elements are masked and grouped into 10 equally sized bins based on the fit of the masked task to the (unmasked) task requirements of a job (see Methods).
-The plot displays the estimated probability that tasks in a bin are required in the job.
-c.
-Predictions of wage offers.
-Jobs are grouped into equally sized bins based on the average value of required tasks.
-The vertical axis shows the estimated mean of the advertised wage offers.
-Vertical bars in panels b and c represent 95% confidence intervals.
-The task space strongly predicts which tasks appear together in job ads.
-To demonstrate this, we combine all job-task vectors into a single dataset and randomly mask 40% of the entries.
-For each masked task, we then measure its relatedness to the remaining unmasked tasks for that job.
-As shown in Fig.
-2b, the probability that a masked task is indeed required (i.e., has a “1” in the task vector) increases sharply with its relatedness to the other tasks, rising 70-fold from below 1% for unrelated tasks to over 12% for strongly related ones.
-The tasks we identify also predict wage offers (Fig.
-2c).
-Salaries in job ads increase alongside the average value of the tasks a job requires, and regression analysis confirms this: a 10% rise in the average value of these tasks is associated with a 9% increase in the offered wage (see Appendix C.2).
-2.3 Task dynamics: growth, diversification, feedback
-Figure 3: Task dynamics.
-a.
-Task user-share change from 2009 to 2022.
-Purple markers signal increases, orange markers decreases, in user-shares between 2009 and 2022.
-Marker transparency reflects the size of shifts: darker tones indicate larger changes in user shares.
-b.
-Estimated probability of diversifying into new tasks at different values of density, users’ relatedness-weighted experience in other tasks: $d_{\theta,u}=\sum_{\kappa}\frac{\mathrm{R}_{\theta,\kappa}}{\sum_{\tau}\mathrm{R}_{\theta,\tau}}X_{\kappa,u}$, where $X_{\kappa,u}$ denotes user $u$’s experience in task $\kappa$ and $\mathrm{R}_{\theta,\kappa}$ the relatedness between tasks $\theta$ and $\kappa$.
-c.
-Regression analysis of answer popularity on task experience.
-Popularity is measured as the number of votes the answer receives, $\log(\#\text{votes}_{a}+1)$ or whether or not the answer is the top answer to the question, task experience as the number of prior answers the user provided to questions on task $\theta$ in the preceding two years, $\log(\#\text{answers}_{u(a),\theta,t(a)})$.
-Control variables include $\log(\#\text{answers}_{q(a)})$ and $\log\left(\sum_{\alpha\in A_{q}}\#\text{votes}_{\alpha}+1\right)$, the total number of answers provided to question $q(a))$, and the sum of all votes across these answers.
-To avoid problems due to $\log 0$ values, we add 1 to counts that can evaluate to 0.
-The plot shows point estimates with their 95% confidence intervals.
-Purple markers refer regression analyses of whether or not an answer is the top answer, orange markers to analyses of the number of votes.
-d.
-Regression of whether or not a user will adopt a new task $\theta$, on the task value, $V_{\theta}$, controlling for the user’s density around the task, $d_{\theta,u}$.
-Purple markers refer to Python-related questions only, orange markers to the full sample.
-Figure 3a illustrates how SO activity shifted across software tasks between 2009 and 2022.
-Notable growth appears in Web Frameworks, certain Android-related tasks, Cloud Computing, and AI/ML.
-Conversely, tasks associated with Advanced Programming Concepts (e.g., Develop a data structure library for efficient data manipulation, Develop a multimedia application in Windows, or Develop a thread-safe application to prevent race conditions), as well as Networking, Enterprise and Web App Development, have seen their relative importance on SO decline.
-We can also apply task vectors to study changes at the individual level, studying how individual SO users develop expertise over time.
-To show this, we generate a task vector for each user by counting how many answers they contribute in a two-year interval for each of the 237 defined tasks (see Methods).
-This approach captures the intensity of engagement in different areas of software work.
-We use these user task vectors to examine how developers add new skills to their profiles over time.
-In particular, we test whether task relatedness—reflected in the task space—predicts which new tasks a user will adopt on SO.
-As shown in Fig.
-3b, users are far more likely to pick up tasks closely related to those they already know, with the probability of transitioning to a “new” task rising from 0.9% (when the task is unrelated) to 14.3% (when it is highly related).
-We also expect that a user’s task-specific experience should correlate with the popularity of their answers.
-Testing this hypothesis is complicated, since users often decide to answer a question only after gauging existing solutions—a process that introduces selection bias.
-In Appendix D.1, we address these biases using instrumental variables.
-Here, we mitigate them partially by focusing on only the earliest answer posted to each question.
-We then ask whether users with greater experience in tasks connected to the question receive more votes.
-Concretely, we regress a binary outcome—whether the answer is the highest-voted—on the user’s prior task experience, controlling for the question’s total number of answers and sum of votes.
-As a robustness check, we also predict the total votes an answer receives.
-Figure 3c reveals a statistically significant, positive link between users’ task experience and the popularity of their answers.
-The estimates indicate that a 10% rise in task experience boosts the likelihood of an answer becoming the top-voted solution by 0.1 pp (±$\pm 0.00004$) and increases the number of votes it receives by 0.14% (±$\pm 0.006\%$).
-In the SI, section D.2 we show that when with sample-selection corrections, the estimated effects grow to 0.006 pp (±$\pm 0.0035$) and 0.9% (±$\pm 0.41\%$), respectively.
-Users also need motivation to adopt new tasks, whether for higher potential rewards or because the tasks align closely with their existing skills.
-To examine these motivations, we regress task entry on the estimated task value (from the aforementioned survey) and the user’s density around the task, i.e., the relatedness-weighted average experience in other tasks the user engages with (see Methods).
-As shown in Fig.
-3d, developers do not universally pursue high-value tasks—likely due to the higher technological barriers or learning costs associated with them—and tend to choose tasks that are more closely related to their current skill sets (see Appendix D.3).
-However, Python users follow a different pattern: task value has a positive effect on their decision to adopt a new skill, suggesting that Python’s versatility allows them to more readily acquire profitable tasks.
-This highlights Python’s distinct role in the software ecosystem: as a general-purpose language, it enables users to branch into new, higher-value tasks more easily, altering and expanding their career trajectories.
-2.4 Tasks and programming languages
-Programmers can draw on a wide variety of programming languages, each with different levels of suitability and popularity for specific software tasks.
-Our observation that Python users often diversify into higher-value tasks suggests that language choice significantly influences developers’ career paths.
-Yet debates persist about which languages best serve which tasks.
-Our task constructs promise a novel way to study language usage.
-To explore this, we use tags that refer to specific programming languages to associate SO posts with the languages they refer to (see Methods).
-This allows us to track which tasks a language supports.
-Figure 4a presents a matrix illustrating which programming languages (used by at least 10 SO users) are used in each of the 237 SO tasks.
-The matrix’s triangular shape signals task nestedness, with an NODF value of 81.5 (Payrató-Borràs et al., 2020), indicating that tasks can be roughly ranked from general to specialized based on how widespread they are across languages.
-Moreover, although different languages may be developed with specific use cases in mind, in practice, this has not led to a strict specialization where tasks are divided between dedicated programming languages.
-Instead, the most specialized tasks are typically handled by languages that cover almost the full spectrum of tasks, while languages used for only a few tasks are largely confined to the most ubiquitous ones.
-Figure 4: Programming languages.
-a.
-Task-language matrix.
-Elements are colored when at least 10 SO users have at least one answer post in the task-language combination.
-b.
-Number of tasks in which a programming language ranks as the top language in terms of SO users.
-The graph shows time-series for the largest six languages in terms of cumulative SO posts between August 2008 and June 2023.
-c.
-Python’s task footprint.
-Nodes are colored if Python ranks among the top 3 programming languages for the associated task.
-Fully colored: 1st rank, 50% transparency: 2nd rank, 75% transparency: 3rd rank.
-Figure 4b illustrates how different programming languages compete for tasks over time.
-Focusing on the six largest languages on Stack Overflow, it tracks how many tasks each language dominates by user counts.
-The most prominent trend is Python’s steady climb: from just nine tasks in 2009, it rises to become the top language for over 80 of the 237 tasks by 2022, surpassing well-established options like C# and JavaScript.
-In the SI, section F, we further demonstrate how languages compete over tasks by focusing on the abrupt shift in iOS programming from Objective-C to Swift.
-Figure 4c visualizes Python’s growth, highlighting all tasks in which it ranks among the top three languages.
-Initially focused on data science, Python spread between 2014 and 2018 into tasks ranging from app development to web-related work, DevOps, and cloud computing.
-Today, it is the dominant language for at least one task in every cluster.
-As detailed in Appendix E, this pattern is preserved when we re-weight languages by their GitHub script counts, a proxy for real-world usage—indicating that Python’s reach extends well beyond Stack Overflow.
-3 Discussion
-The growing complexity of the labor market has made measuring worker capabilities both more difficult and more important (Anderson, 2017).
-Here we demonstrated that data on individual problem solving can be used to generate a useful fine-grained taxonomy of tasks people do in software-related work, a key industry in the digital economy.
-These tasks or building blocks of software work can be given coherent economic values and capture co-occurrence relationships reflected in real world job advertisements.
-We use the tasks to observe individual learning, both within tasks and entrance into new ones.
-Moreover, the fine-grained detail of the task space allows us to explain an important macro-system outcome: the rise of Python as the leading all-purpose programming language.
-Our explanation for Python’s success is that it enables people to enter new tasks which have higher economic value.
-Ultimately, our results highlight how deepening our understanding of the microstructure of tasks offers new ways to interpret macro-level phenomena such as technology diffusion, wage dispersion, and the evolution of industry-wide practices.
-A core insight from our results is that tasks exist in a space of related capabilities, echoing network-based accounts of human capital (Anderson, 2017).
-On the one hand, tasks are well-defined as distinctive categories of specific kinds of work, as evidenced by their distinct labor market values and that effective learning happens by practice within tasks (Arrow, 1962).
-On the other hand, tasks also have meaningful relationships with each other, evidenced by their non-random co-occurrence within job ads and observed patterns of individuals learning new tasks.
-The resulting networked taxonomy of software tasks demonstrates its value by providing a micro-to-macro explanation for the rise of the Python programming language: Python enables users to enter more valuable tasks.
-Our findings connect with the literature on skill complexity, labor market transitions, and the organization of work, complementing studies that have emphasized how granular task analysis can reveal hidden dynamics in career mobility, wage disparities, and skill complementarities (Autor et al., 2003; Anderson, 2017; Alabdulkareem et al., 2018; del Rio-Chanona et al., 2021).
-Previous research however, tends to use standardized task classifications such as O*NET to define tasks and applies them to curated job descriptions, limiting the granularity and responsiveness of task taxonomies.
-In contrast, by extracting tasks directly from real-world problem-solving interactions on Stack Overflow, our approach captures the immediate and practical realities of skill application at a global scale.
-In rapidly changing sectors such as software engineering, tasks derived from activities can explain surprising developments like the rise of the Python programming language as a technology which enables diversification into more valuable tasks.
-This complements existing work on task-relatedness and collective learning.
-We highlight three possible applications of our method and taxonomy.
-First we have demonstrated that microdata on activity can be fruitfully aggregated into tasks to describe skill acquisition, development and transitions.
-Tasks from any sector in which large-scale fine-grained activity on individual behavior at work is available could be categorized in a similar manner.
-Second, employers in such sectors can use our approach to identify skills related to tasks their employees actually carry out, enabling better labor market matching outcomes or internal skill development programs.
-Finally, educators can better understand the labor market landscape of the software industry using our taxonomy.
-There are several limitations to our study, which may be tackled in future work.
-First, we have focused exclusively on programming-related tasks.
-However, software developers need to be able to carry out many other tasks, such as managing teams or reporting results.
-Although SO contains some postings about organizational aspects of software development, its emphasis is on technical subjects.
-To explore how programming tasks are combined with other types of tasks in software jobs, future studies could use job ads to compare the co-occurrence of our programming tasks with tasks related to different kinds of soft skills.
-Second, although the question-answer structure makes SO ideal for identifying software tasks, it may be less suited for descriptions of macro-level trends and individual career paths.
-In Appendix E, we compare user numbers on SO to GitHub.
-This analysis suggests that in terms of capturing software developers, both datasets are closely aligned.
-However, it is impossible to judge the representativity of SO of overall software activity at the level of tasks, especially as Stack Overflow is thought to be declining (del Rio-Chanona et al., 2024).
-In Appendix E, we therefore rescale the activity in different programming languages such that they reflect their presence on GitHub.
-However, in future work, we plan to define software tasks directly in GitHub.
-Third, we opted to create a task taxonomy that is time-invariant.
-However, an analysis of how job tasks change in software development would provide a more vivid picture of the software industry.
-Such an analysis could also expand the size of our job ad dataset.
-Although sufficient for the validation exercise we conducted, a larger dataset would allow us to see how tasks in software development and their value change over time.
-Notwithstanding these limitations, our study demonstrates how question-and-answer data can inform us about the evolution of work by quantifying the abstract notion of a task.
-The application to the software sector sheds light on novel aspects of high skill labor markets at high granularity, at scale, and in near real-time.
-Moreover, programming tasks are rapidly diffusing beyond pure software development firms and understanding these tasks and their evolution is pivotal to understanding the future of work in an increasingly digital economy.
-4 Methods
+These challenges reflect real-world coding tasks and by posting answers users demonstrate skills and expertise in these tasks.
+Indeed, SO postings have real-world value as signals to prospective employers of a programmer’s capabilities (Xu et al., 2020).
+Interpreting tags as a structured, albeit partial, description of posted questions, we aggregate questions to create a software-skill taxonomy by identifying clusters of frequently co-occurring tags.
+We then introduce a notion of relatedness among programming skills based on co-occurrence patterns in users’ posting histories.
+We use this relatedness to construct a skill space where skills are connected if they are often jointly possessed by the same users.
+We then use this taxonomy and skill space to examine the postings and learning trajectories of programmers.
+We find that our skill constructs can be used to predict salaries in real-world job postings.
+Notably, wages vary greatly across skill profiles, with higher-paying jobs concentrated in more specialized or in-demand fields such as AI and machine learning.
+Similarly, relatedness estimates derived from interactions with the SO platform predict which combinations of skills are demanded in real-world job ads.
+This not only shows that our taxonomy captures meaningful structures in the software development landscape, but also that software development jobs are coherent in the sense that they often require sets of closely related skills.
+The skill taxonomy also sheds light on the role and evolution of programming languages.
+Programming languages differ in the expertise they support.
+Analyzing which skills are used with which programming languages, we observe a “nested” pattern reminiscent of similar patterns observed in ecology (Bascompte et al., 2003), economic development (Hidalgo and Hausmann, 2009; Tacchella et al., 2012; Mariani et al., 2019) and skill hierarchies (Hosseinioun et al., 2025).
+This means that although programming languages may have been tailored for specific use cases, this is not observed in how programmers use languages in practice.
+Instead, we observe that skills have an implicit hierarchy such that the most ubiquitous skills are supported by almost all languages, whereas rare skills often rely on the most versatile ones.
+A prime example of the latter is the programming language Python, a popular language that has seen significant growth in recent years (Economist, 2018).
+Our skill taxonomy shows that this growth has been accompanied by changes in the way Python is used, with Python becoming increasingly versatile, rapidly increasing the variety of skills it supports to become today’s premier general purpose language.
+At the individual user level, we observe two patterns.
+First, developers excel —as inferred from community feedback on their posts— at the skills they are most active in.
+Second, developers acquire new skills over their SO careers, and the relatedness network strongly influences which skills they learn: developers are fifteen times more likely to acquire skills related to skills they already master than to unrelated ones.
+At the same time, we find that developers usually add skills with lower market value, which are presumably easier to learn.
+However, Python enables qualitatively different learning paths.
+Specifically, unlike users of most other languages, Python users often branch into more, not less valuable skills.
+This observation could help explain Python’s success, despite its not being the oldest or fastest language: its ecosystem may make it easier for developers to develop higher-value expertise.
+Our analysis contributes to the growing literature that analyzes the content of work.
+This literature has shown that as wages and careers have come to depend on how workers specialize (Hosseinioun et al., 2025), interact (Deming, 2017), complement one another (Neffke, 2019), and move across places and jobs (Frank et al., 2024), fine-grained skill taxonomies are increasingly valuable.
+Such taxonomies clarify how workers transition between occupations (del Rio-Chanona et al., 2021), how skills substitute for or complement each other (Anderson et al., 2012; Neffke, 2019), and how occupations evolve over time (Nedelkoska et al., 2021).
+This line of research uses data on skill requirements and work activities extracted from large-scale surveys (O*NET, 2025), expert assessments (Commission, 2017), job ads (Börner et al., 2018) or online freelance platforms(Anderson, 2017; Stephany and Teutloff, 2024) to describe jobs as networks of skills or tasks.
+We contribute to this debate by focusing away from the task requirements listed in descriptions of jobs toward the human capital and skills of the people who fulfill them.
+This allows us to study how workers specialize and learn new things, processes critical to a successful career.
+2 Mapping software work
+In this section, we describe how we construct a software skill taxonomy and a space of software skills.
+First we describe the raw inputs: posts on Stack Overflow and their tags, labels which capture post content.
+We then explain how we use a network-based clustering algorithm, the stochastic block model (SBM, see Appendix A), to identify clusters of tags which frequently co-occur in posts.
+These clusters are our skills, from which we can derive descriptions of user skills and skill values.
+Finally we demonstrate how skills can be embedded in an abstract space, in which related skills are near one another.
 2.1 Data
-Stack Overflow Posts:
 Our main dataset is the corpus of questions and answers (together “posts”) on Stack Overflow from the site’s launch in August 2008 to June 2023.
@@
 For the latter, we merge different versions of the same programming language, converting, for instance python-3.x into the generic tag python.
-When we define tasks, we use all available posts, with answers inheriting the tags of the questions they answer.
-In all other analyses, we only consider answers from users with a total of at least 10 answer posts.
-Stack Overflow Develop Survey:
-Stack Overflow provides access to yearly surveys of its users, the Stack Overflow Developer Survey, with anonymous information on salaries, as well as on tools and languages used that can often be linked to tags in SO.
-We use the survey data from year 2023, which contains around 12,000 participants who report their salaries to estimate task values.
-HackerNews Job Ads:
-We scraped job advertisements from “hnhiring.com”, an index of jobs from the website Hacker News’ “Who is Hiring?” posts.
-This yielded a total of 50,998 job advertisements.
-We use approximately 46,000 ads to study task co-occurrence and 3,949 jobs to test the ability of tasks to predict salaries.
-4.2 Building software tasks
-4.2.1 Identifying programming languages
+When we define skills, we use all available posts, with answers inheriting the tags of the questions they answer.
+When we analyze users, we only use their answer posts (ignoring the questions they post) to assess the content of their human capital.
+To ensure sufficiently accurate descriptions of this human capital, we restrict our analysis to users who post at least 10 answers over the entire period our data cover.
+Stack Overflow also provides access to yearly surveys of its users, the Stack Overflow Developer Survey.
+This provides anonymous, self-reported information on salaries, as well as on tools and languages used that can often be linked to tags in SO.
+We use the survey data from year 2023, which contains around 12,000 U.S.
+participants who report salary information.
+Finally, we scraped job advertisements from “hnhiring.com”, an index of jobs from the website [^3] https://www.hackernews.com/ Hacker News’ “Who is Hiring?” posts.
+This yielded a total of 50,998 software related job advertisements.
+We use approximately 46,000 ads to study skill requirements of jobs and 3,949 ads listing salary information to test the ability of skill requirements to predict wage offers.
+2.2 Constructing a taxonomy of software skills
+Our goal is to construct a taxonomy of skills related to tasks in software development work.
+Software development tasks can be viewed as recurring categories of programming problems that developers encounter.
+The ability to provide advice on how to tackle such problems demonstrates skills and expertise in the area.
+Therefore, SO, with its vast repository of user-generated questions and answers, offers a unique window onto these real-world challenges and the human capital required to meet them.
+In order to organize its community and facilitate information retrieval, the SO platform labels each question with one or more of 66,000 curated tags that indicate the specific tools, frameworks, or concepts involved.
+Tags are deduplicated and moderated by automated bots and human moderators, ensuring relatively high quality.
+Before constructing a skill taxonomy, we filter tags in two ways.
+First, we drop tags appearing fewer than 1000 times.
+Second, we distinguish tags that refer to programming languages rather than concepts or methods.
+We drop these tags from our skill taxonomy, because we programming languages are rather tools to implement software skills; the same skill could be applied in different languages.
+While languages themselves are specialized (Valverde and Solé, 2015; Juhász et al., 2026), they are a different category.
+In fact, we will later analyze the relationship between skills and languages used.
 To identify tags that refer to programming languages, we rely on wiki pages, which are available for most common tags.
 We first filter the pages to find tags that include “language” or “framework” among their keywords.
-We then manually review all the tags to select those that refer to programming languages, yielding a total of 282 programming languages.
+We then manually review these tags to select those that refer to programming languages, yielding a total of 282 programming languages.
 Some programming tags refer to specific versions of a language, such as “python-3.x” or “python-3.7”.
-When calculating the nestedness of the task-language matrix, we merge all tags referring to different versions of the same programming language.
-Yet, there is some ambiguity to what counts as a programming language, such as frameworks and developer environments.
-We remove such tags when calculating nestedness, basing this estimate on a task-language matrix with 114 programming languages.
-4.2.2 Identifying software tasks
-To identify canonical software tasks, we express the relation between questions and their tags as a bipartite network with 18,154,593 question nodes that connect to 5,083 tags through 33,378,590 edges.
-We then apply a bipartite Stochastic Block Model (SBM) to identify communities of tags.
+In these cases, we merge all tags referring to different versions of the same programming language.
+Yet, there is some ambiguity as to what counts as a programming language (e.g., web frameworks and developer environments, which are not fully fledged programming languages).
+We remove such tags when analyzing the relation between skills and languages, yielding 114 unambiguous instances of programming languages.
+2.2.1 Identifying software skills
+With the processed set of tags in hand, we can now identify canonical software skills in our data set.
+To do so we express the relation between questions and their tags as a bipartite network with 18,154,593 question nodes that connect to 5,083 tags through 33,378,590 edges.
+We use these tag–question relationships to group related tags into “canonical” areas of software development expertise (“skills”), applying an SBM (Peixoto, 2017, 2014; Gerlach et al., 2018) to detect communities in the bipartite graph that connects questions to tags (Fig.
+1a).
 SBMs essentially estimate a coarsened version of a network.
-To do so, they identify communities of nodes and parameters that represent parsimonious descriptions of edge distributions within and between communities.
-To sharpen the boundaries of inferred communities, we remove $20\%$ of tags that have the weakest own-community attachment (see Appendix section A).
-Dropping communities with fewer than 3 tags, we retain 237 tasks, connected to 4,054 tags, out of originally 247 detected communities.
-4.3 User task vectors
-We use the tasks extracted from the question-tag network to describe the activity of individual SO users as a 237-dimensional vector.
-Entries of this vector count how many answers the user has provided to questions that are labeled with tags that are associated with each corresponding task.
-Specifically, we express the user’s task experience in time period $t$ as:
-$\vec{T}_{u}^{t}=\begin{bmatrix}X_{u,1}^{t}\\ ...\\ X_{u,\theta}^{t}\\ ...\\ X_{u,N_{\theta}}^{t}\\ \end{bmatrix},$ (1)
-where $N_{\theta}=237$ the total number of tasks and $X_{u,\theta}^{t}$ is the number of answers posted in period $t$ by user $u$ to questions labeled with a tag belonging to the community of tags associated with task $\theta$.
-4.4 Constructing a Task Space
-To calculate the relatedness between tasks, we analyze which tasks are often associated with the same users.
-To do so, we collect the task vectors of all users in an $(N_{\theta}\times N_{u})$ matrix T, with $N_{u}$ the number of users in the dataset with at least 10 answers and $N_{\theta}=237$ the total number of tasks.
-Next, we ask how often two tasks co-occur within the same users by multiplying $(\textbf{T}>0)$ with its transpose: $(\textbf{T}>0)(\textbf{T}>0)^{\prime}$.
-Finally, we calculate the Pointwise Mutual Information ($\mathrm{PMI}$) of probability $p_{\theta,\kappa}$ that task $\theta$ and $\kappa$ co-occur in the same user to capture the amount of (information-theoretic) surprise involved in observing the joint probability $p_{\theta,\kappa}$ against a benchmark where draws of tasks $\theta$ and $\kappa$ are independent:
-$\mathrm{PMI}_{\theta,\kappa}=\log\frac{p_{\theta,\kappa}}{p_{\theta}p_{\kappa}},$ (2)
+To do so, they identify communities of nodes and estimate parameters that yield parsimonious descriptions of edge distributions within and between communities.
+To sharpen the boundaries of inferred communities, we remove $20\%$ of tags that are only weakly connected to the other tags in their community (see Appendix A).
+This yields 247 different communities.
+Dropping communities with fewer than 3 tags, we retain 237 communities, connected to 4,054 tags, which we will henceforth will refer to as “skills”.
+To generate concise names for each skill community, we label them by an example of a programming skill that illustrates the tags in the community.
+That is, we name each skill after the task it allows to perform.
+To produce these labels, we prompt the large language model ChatPGT-4.0 as follows:
+I have a list of tags related to software engineering and computer science.
+The tags are ordered by their importance, listing the most important task first.
+Using no more than 10 words, could you please suggest a general task related to the following tags ’task_description’ as key words?
+Please make sure that the task is general and not too specific.
+Please only return json content and not anything else.
+This process yields labels that capture the essence of the tags comprising each group (see Appendix B), such as develop NLP models and tools for text analysis, establish secure server access and control, and deploy and secure a scalable web app.
+2.3 User skill vectors
+Skills are rarely valuable in isolation (Neffke, 2019).
+Therefore, jobs often require multiple skills.
+When skills complement each other, they are often required in the same jobs (Anderson, 2017; Alabdulkareem et al., 2018; Stephany and Teutloff, 2024).
+We investigate this skill bundling by identifying which combinations of skills are frequently mastered by the same SO users.
+To assess to what extent a user masters a skill, we rely only on the answers they post, not on the questions they raise.
+We use the skills (defined from the tasks they allow performing) extracted from the question-tag network to describe the activity of individual SO users as a 237-dimensional vector.
+Entries of this vector count how many answers the user has provided to questions that are labeled with tags that are associated with each corresponding skill.
+Specifically, we express the user’s expertise in time period $t$ as:
+$\vec{S}_{u}^{t}=\begin{bmatrix}X_{u,1}^{t}\\ ...\\ X_{u,\theta}^{t}\\ ...\\ X_{u,N_{\theta}}^{t}\\ \end{bmatrix}$, (1)
+where $N_{\theta}=237$ is the total number of skills in our taxonomy and $X_{u,\theta}^{t}$ the number of answers posted in period $t$ by user $u$ to questions labeled with a tag belonging to the community of tags associated with skill $\theta$.
+We will interpret this number as the user’s expertise intensity in a skill.
+2.4 Skill values
+We estimate the value of a skill based on salary information from the Stack Overflow developers survey.
+Because wages differ widely across countries, we focus on 12,000 US respondents that report a salary in 2023.
+Respondents list the technologies they use, which often correspond to tags in the SO platform.
+We use tags that are used in both the survey and the general SO database to find closely matching survey respondents for each SO user.
+Selecting the 300 most similar survey respondents, we calculate the weighted average salary reported by these respondents.
+We will use this average as a rough estimate of the wage that the user could command on the US labor market:
+$V_{u}$ $=\sum_{\rho\in R_{u}}\frac{M_{u,\rho}}{\sum_{\alpha}M_{u,\alpha}}V_{\rho},$ (2)
+where $V_{u}$ is user $u$’s inferred US wage, $V_{\rho}$ the wage of survey respondent $\rho$, and $R_{u}$ the set of survey respondents we matched to user $u$.
+Weights, $M_{u,\rho}$, are given by the number of tags that user $u$ and respondent $\rho$ share.
+To assign a value to a skill, we next calculate the weighted average salary of users possessing the skill, using the skill’s intensity in users’ skill vectors for the years 2018-2023 (the period in which we collect job ads) as weights:
+$V_{\theta}$ $=\sum_{u}\frac{X_{u,\theta}^{2018,2023}}{\sum_{u^{\prime}}X_{u^{\prime},\theta}^{2018,2023}}V_{u}.$ (3)
+2.5 Constructing a skill space
+To determine the relatedness between skills, we analyze which skills are often possessed by the same users.
+To do so, we collect the skill vectors of all users in an $(N_{\theta}\times N_{u})$ matrix S, with $N_{u}$ the total number of users in the dataset and $N_{\theta}=237$ the total number of skills.
+Next, we ask how often two skills co-occur within the same users by multiplying $1(\textbf{S}>0)$ with its transpose: $1(\textbf{S}>0)1(\textbf{S}>0)^{\prime}$, where $1(.)$ is an indicator function that evaluates to 1 if its argument is true.
+Finally, we calculate the Pointwise Mutual Information ($\mathrm{PMI}$) of probability $p_{\theta,\kappa}$ that skill $\theta$ and $\kappa$ co-occur in the same user to capture the amount of (information-theoretic) surprise involved in observing the joint probability $p_{\theta,\kappa}$ against a benchmark where draws of skills $\theta$ and $\kappa$ are independent:
+$\mathrm{R}_{\theta,\kappa}=\mathrm{PMI}_{\theta,\kappa}=\log\frac{p_{\theta,\kappa}}{p_{\theta}p_{\kappa}},$ (4)
 where $p_{\theta}$ and $p_{\kappa}$ are marginal probabilities, i.e., $p_{\theta}=\sum_{\gamma}p_{\theta,\gamma}$ and $p_{\kappa}=\sum_{\mu}p_{\mu,\kappa}$.
-To estimate $\mathrm{PMI}$ values and their credible intervals, we apply a Bayesian statistical framework developed by (van Dam et al., 2023).
-The task space is based on a weighted network where nodes represent tasks and edges the relatedness between them.
-Because our focus is on relatedness as expressed in a surprisingly large observed probability that two tasks co-occur, we ignore tasks that co-occur less frequently than our random null model predicts by setting all negative $\mathrm{PMI}$ values to 0.
-To identify clusters in the task space, we embed the relatedness network in a 5-dimensional space using UMAP (McInnes et al., 2018).
-Next, we cluster sets of contiguous tasks using HDBSCAN (McInnes et al., 2017).
+We estimate $\mathrm{PMI}$ values and their credible intervals using a Bayesian statistical framework developed by (van Dam et al., 2023).
+We will refer to these values as the relatedness between skills.
+We use these relatedness estimates to construct a skill space, a network where nodes represent skills connected by edges that indicate the relatedness between them.
+In doing so, we only connect skills with positive relatedness values, ignoring negative relatedness (where skills co-occur less frequently than our random null model predicts) by setting all negative $\mathrm{PMI}$ values to 0.
+To help interpret this skill space, we embed the relatedness network in a 5-dimensional space using UMAP (McInnes et al., 2018).
+Next, we cluster sets of contiguous skills using HDBSCAN (McInnes et al., 2017).
 These clusters are represented in colors of Fig.
 1b.
-The nodes’ coordinates in this figure are determined by a UMAP embedding of the relatedness network in a 2-dimensional space.
-4.5 Task Values
-We estimate the value of a task based on salary information from the Stack Overflow developers survey.
-We focus on the 12,000 US participants that report a salary in 2023.
-Respondents list the technologies they use, which often correspond to tags in the SO platform.
-For each SO user, we use the top 300 most similar survey respondents in terms of overlap in tags and technologies to estimate user salary.
-We use the weighted average salary reported by the most similar respondents, weighting by tag similarity to the SO user.
-To assign values to tasks, we then calculate the weighted average salary of the users that carry out that task, using the prevalence of the task in users’ task vectors for the years 2018-2023 (the period in which we collect job ads) as weights:
-$V_{\theta}$ $=\sum_{u}\frac{X_{u,\theta}^{2018,2023}}{\sum_{u^{\prime}}X_{u^{\prime},\theta}^{2018,2023}}V_{u}$ (3)
-$V_{u}$ $=\sum_{\rho\in R_{u}}\frac{M_{u,\rho}}{\sum_{\alpha}M_{u,\alpha}}V_{\rho},$ (4)
-where $M_{u,\rho}$ is the number of overlapping tags between SO user $u$ and one of the 300 closest survey respondents, $\rho\in R_{u}$, and $V_{\rho}$ is the salary of survey respondent $\rho$, such that $V_{u}$, the match-weighted average salary of survey respondents matched to user $u$, can be interpreted as a proxy for the user’s salary.
-4.6 Job Advertisements
-The data on job ads come from the website HackerNews, and were scraped from the aggregator site hnhiring.com.
-They represent all 50,998 job ads posted to the website between January 2018 and May 2024.
-These ads are unstructured job descriptions sometimes including salary and expected work activities (tasks).
-We prompted an LLM (ChatGPT-3.5) to extract salary information and task requirements, which yields about 46,000 jobs listing at least 3 tasks, and 3,949 jobs that contain at least 1 task and salary information.
+The node coordinates in this figure are determined by a UMAP embedding of the relatedness network in a 2-dimensional space.
+Note, however, that neither the clusters nor the visual representation of the skill space will play a role in our quantitative analyses.
+Clusters often group skills with comparable labor market returns.
+This is illustrated in Fig.
+1d, where, we plot the inferred skill values of eq.
+(3).
+For example, on the right side of the skill space, we find three clusters revolving around website development that are typically linked to lower wages.
+By contrast, skills in the top center—focusing on advanced iOS and Android development—offer significantly higher earnings.
+Additional high-value clusters include those for AI and machine learning (AI/ML), Advanced Programming Concepts, Cloud Computing, and DevOps.
+The table of Fig.
+1e lists the five highest-paying skills, three of which are related to iOS app development, and the five lowest-paying skills, all associated with more basic web development.
+Notably, clusters that are conceptually linked—such as AI/ML and Statistics and Data Analysis, and iOS and Android development—tend to lie near each other in the skill space.
+To further investigate the network structure of the skill space, we can examine individual skills in greater detail.
+For example, the ten skills most closely related to developing AI models (Fig.
+1c) reveal strong connections to other skills in statistics, as well as skills related to image and natural language processing, data visualization, and packaging models into standalone applications.
+Fig.
+B.2 shows changes in the relative importance of different skills, highlighting important areas of growth around skills in web frameworks, but also in high-value skills related to AI, cloud computing and Android.
+Figure 1: Mapping software skills.
+a.
+Stylized depiction of the bipartite question-tag network.
+SBM groups tags into communities (skills) that connect to similar sets of questions.
+ChatGPT-4.0 finds a common label that summarizes each community’s tag information.
+b.
+Skill space.
+Pointwise mutual information (relatedness) expresses how surprisingly often two skills are possessed by the same users.
+UMAP embeds the resulting relatedness network in a 2-dimensional plane (the skill space).
+c.
+Close-up on the skill to develop AI models, depicting the original network structure among the 10 most closely related skills.
+d.
+Skill values.
+Nodes are colored according to their skill value, estimated from salary information in the 2023 SO developers survey.
+Darker shades indicate more valuable skills.
+e.
+Table of the five most and least valuable skills.
+3 Results
+3.1 Validation in job ads
+How well do these software skills capture real-world development work beyond SO?
+To find out, we examine all 50,998 job ads posted to the Hacker News website between January 2018 and May 2024.
+These ads are unstructured job descriptions, sometimes including salary and expected work activities (tasks).
+We prompt an LLM (ChatGPT-3.5) to extract salary information and work activities (see Appendix C.1), which yields about 46,000 jobs listing at least 3 work activities, and 3,949 jobs that contain at least 1 work activity as well as salary information.
 When ads describe more than one position, the LLM provides information on each job separately.
-We first relate the task requirements of job ads to the 237 SO tasks.
-We embed both job ad tasks and SO tags in a 384-dimensional vector space using sBERT (Reimers and Gurevych, 2019).
-To describe an SO task in this vector space, we calculate the embedding of its label.
-To increase accuracy, we use somewhat longer labels of up to 30 words generated by ChatGPT-4.0.
-Finally, we match each task requirement of a job ad to the closest SO tasks based on cosine similarity between the embeddings of the job-ad task and of the SO task.
-When we cannot find any SO task with a cosine similarity of at least 0.3, we drop the task requirements.
-This yields a 237-dimensional vector of zeros and ones that describe each job’s requirements in terms of our SO tasks.
-In Appendix C.3, we demonstrate robustness of our findings to alternative methods and thresholds.
-We use these data in two ways to test the validity of our SO tasks.
-First, we ask whether we can predict the task composition of a job.
-To do so, we take the SO-task vectors associated with each job ad and stack these vectors into a dataset that contains for each combination of an SO task and an advertised job whether or not the task is required by the job.
-Next, we mask 40% of observations in this dataset and calculate the density of required tasks in the job around the masked tasks:
-$\mathrm{D}_{\theta,j}=\bar{\mathrm{R}}_{\theta,j}\mathrm{T}_{j},$ (5)
-where $\bar{}$ denotes row-normalization and $\mathrm{T_{j}}$ and $\mathrm{R}_{\theta,j}$ are the unmasked portions of job $j$’s task vector and the corresponding rows of relatedness matrix $\mathrm{R}$, namely $\mathrm{R}_{\theta,i}=\frac{\mathrm{R}_{\theta,i}}{\sum_{\tau}\mathrm{R}_{\theta,\tau}}$.
-$\mathrm{D}_{\theta,j}$ denotes the average relatedness of (masked) task $\theta$ to all unmasked tasks in job $j$.
-We plot how the estimated probability that a job requires a particular (masked) task changes with that task’s density in the job’s unmasked task requirements.
-To do so, we bin all masked observations in 10 equally sized bins and calculate relative frequencies with their 95% confidence intervals.
-These estimated probabilities are plotted on the vertical axis of Fig.
-2b, against the bin’s average task density on the horizontal axis.
-Our second validation focuses on the set of advertised jobs that also post salary information.
-It estimates the association between the average value of a job’s SO tasks and the posted wage in the job ad.
-We report on this in SI, section C.2.
-4.7 Analysis of SO users
-To study user behavior, we randomly divide the population of users into two equally sized samples.
-We use the first sample, $S_{1}$, to construct a task relatedness matrix from co-occurrence patterns of tasks in users.
-The second sample, $S_{2}$, is used to estimate how relatedness affects user activity on SO.
-To do so, we collect information from all answer posts by these users within rolling two-year intervals.
-That is, we construct experience vectors $\vec{T}_{u}^{t}$, where $t$ denotes an interval that stretches across the two calendar years preceding year t.
-For instance, $\vec{T}_{u}^{2018}$ refers to the answers user $u$ provides between January 1, 2016 and December 31, 2017.
-4.7.1 Task diversification
-Analogously to eq. (5), we use sample $S_{2}$ to calculate for a given user, $u$, and year $t$, the experience density around each task:
-$\mathrm{D}^{t}_{u}=\bar{\mathrm{R}}_{S_{1}}\mathrm{T}^{t}_{u},$ (6)
-where $\bar{\mathrm{R}}_{S_{1}}$ is the row-normalized relatedness matrix, constructed using users in sample $S_{1}$.
-Next, we ask which new tasks users are most likely to enter.
-To do so, we take all tasks in which user $u$ had no activity in the two calendar years preceding $t$.
-These tasks represent the set of tasks that user $u$ can enter.
-We tag all such tasks with a 1 whenever the user posts at least one answer related the the task in year $t$, and with 0 otherwise.
-Next, we divide all user-task combinations at risk of entry into 10 equally sized bins and estimate the probability that users do enter these tasks in year $t$.
-Figure 3b shows how these estimated probabilities rise with the user’s experience density around a task.
+Next, we match each extracted work activity to the closest skill in our taxonomy.
+To do so, we embed work activities and skills in a 384-dimensional vector space using sBERT (Reimers and Gurevych, 2019).
+For work activities, we embed the phrases extracted by the LLM; for skills, we calculate the embedding of their labels.
+[^4] To increase accuracy, we use somewhat longer labels of up to 30 words generated by ChatGPT-4.0.
+Finally, we match each work activity in a job ad to the closest skill based on the cosine similarity between the embeddings of the work activity and the skills.
+When we cannot find any skill with a cosine similarity of at least 0.3, we drop the work activity.
+This yields a 237-dimensional vector of zeros and ones that describe each job’s requirements in terms of the skills of our taxonomy.
+[^5] In Appendix C.3, we show that our findings are robustness when we change this threshold or when use embeddings for skills based directly on their tags instead of their labels.
+We combine this vector with information on skill-values and relatedness among skills.
+Fig.
+2a illustrates this process.
+We use these data in two ways to test the validity of our skill taxonomy.
+Our first validation focuses on the set of advertised jobs that also post salary information.
+It estimates the association between the average value of a job’s skills and the posted wage in the job ad.
+We find that skill requirements help predict wage offers (Fig.
+2b).
+Salaries in job ads increase alongside the average value of the skills a job requires.
+Regression analysis confirms this: a 10% rise in the average value of these skills is associated with a 9% increase in the wage offer for the advertised job (see Appendix C.2).
+Second, we ask whether we can predict the skill composition of a job.
+To do so, we take the skill vectors associated with each job ad and stack these vectors into a dataset that contains for each combination of a skill and an advertised job whether or not the skill is required in the job.
+Next, we mask 40% of observations in this dataset and calculate the average relatedness of a masked skill requirement, $\theta$, to all unmasked skill requirements in the same job, $j$:
+$\mathrm{D}_{\theta,j}=\bm{\bar{\mathrm{r}}}_{\theta}^{j}\bm{\mathrm{s}}^{j},$ (5)
+where $\bm{\mathrm{s}}^{j}$ is a vector that subsets the unmasked (i.e., “visible”) portions of job $j$’s skill requirement vector, $S_{j}$.
+$\bm{\bar{\mathrm{r}}}_{\theta}^{j}$ refers to the corresponding row-normalized vector of relatedness matrix $\mathrm{R}$, consisting of elements $\frac{\mathrm{R}_{\theta,\zeta}}{\sum_{\tau\in\Xi_{j}}\mathrm{R}_{\theta,\tau}}$, where $\Xi_{j}$, the set of unmasked skills in the skill requirements vector of job $j$, such that $\zeta\in\Xi_{j}$ and $\theta\notin\Xi_{j}$.
+The skill space strongly predicts which skill requirements appear together in job ads.
+To show this, Fig.
+2c plots how the estimated probability that a job requires a particular (masked) skill changes with that skill’s density in the job’s unmasked skill requirements.
+First, we sort all masked observations by the unmasked skill’s density in the job, then bin them in 10 equally sized bins and calculate relative frequencies.
+Next, we plot these estimated probabilities on the vertical axis against the bin’s average skill density on the horizontal axis.
 95% confidence intervals are approximated with a normal distribution, using the following estimate of the standard error of the mean:
 $\hat{\sigma}_{b}=\frac{1}{\sqrt{n_{b}}}\sqrt{\hat{\pi}_{b}(1-\hat{\pi}_{b})},$ (6)
-where $n_{b}$ the number of observations in bin $b$ and $\hat{\pi}=\frac{d_{b}}{n_{b}}$, where $d_{b}$ the number of entered tasks in bin $b$, the relative frequency with which tasks in bin $b$ are entered.
-4.7.2 Voting
-To study of a user’s proficiency and learning in a task, we analyze the votes the user’s answers receive.
+where $n_{b}$ the number of observations in bin $b$ and $\hat{\pi}=\frac{d_{b}}{n_{b}}$ (where $d_{b}$ the number of skills required in jobs in bin $b$) representing the relative frequency with which masked skills are required in bin $b$.
+The probability that a masked skill is required (i.e., has a “1” in the skill requirements vector) increases sharply with its relatedness to the unmasked skill requirements, rising almost 20-fold from 1.0% for unrelated skills to 19.6% for strongly related ones.
+This not only shows that our SO-based skill taxonomy is meaningful in the real-world software development labor market, but also that software development jobs are coherent: they, on average, require skill combinations that are often mastered by the same users on SO.
+Figure 2: Job ads.
+a.
+Schematic representation of the workflow to extract salary and work activities from online job ads by prompting ChatGPT.
+Work activities are converted to the 237-dimensional skill requirement vectors based on cosine similarities between text embeddings of work activities and skills.
+b.
+Prediction of wage offers.
+Jobs are grouped into equally sized bins based on the average value of required skills.
+The vertical axis shows the estimated mean of the advertised wage offers.
+c.
+Prediction of skill requirements.
+40% of elements in each skill requirement vector are masked and grouped into 10 equally sized bins based on the average proximity of a masked skill requirement to all unmasked skill requirements in a job.
+The plot displays the estimated probability that a specific masked skill in a bin is required for a job.
+Vertical bars in panels b and c represent 95% confidence intervals.
+3.2 Mapping skills to programming languages
+Programmers can draw on a wide variety of programming languages, each with different levels of suitability and popularity for specific use cases.
+However, which languages best serve which areas of software development remains hotly debated.
+Our skill constructs promise a novel way to approach this question: using tags that refer to specific programming languages we can associate SO posts with the languages they refer to (see section 2.2).
+This, in turn, allows us to track which skills are used with which languages.
+Fig.
+3a presents a matrix illustrating which programming languages are used by at least 10 users with each of the 237 skills we identified, based on answer postings whose tags refer to both the skill and the language.
+A visual inspection suggests that the matrix is triangular.
+This indicates that skills are nested (Hosseinioun et al., 2025), as confirmed by an NODF value of 81.5 (Payrató-Borràs et al., 2020).
+Accordingly, skills can be roughly ranked from general to specialized based on how widely they are implemented across languages.
+That is, although different languages may be developed with specific use cases in mind, in practice, this has not led to a strict specialization where dedicated programming languages serve specific areas of software development expertise.
+Instead, whereas common skills can be performed in almost any language, highly specialized skills often require generalist languages.
+Figure 3: Programming languages.
+a.
+Skill-language matrix.
+Elements are colored when at least 10 SO users have at least one answer post in the skill-language combination.
+b.
+Number of skills for which a programming language ranks as the top language in terms of SO users.
+The graph shows time-series for the six largest languages in terms of cumulative SO posts between August 2008 and June 2023.
+c.
+Python’s skill footprint.
+Nodes are colored if Python ranks among the top 3 programming languages for the associated skill.
+Fully colored: 1st rank, 50% transparency: 2nd rank, 75% transparency: 3rd rank.
+Fig.
+3b illustrates how different programming languages compete for user types over time.
+Focusing on the six largest languages on Stack Overflow, it tracks how many distinct skill areas each language dominates by user counts.
+The most prominent trend is Python’s steady climb: from just nine skills in 2009, it rises to become the top language for over 80 of the 237 skills by 2022, surpassing well-established options like C# and JavaScript.
+In the Appendix F, we further demonstrate how languages compete for dominance in expertise areas by focusing on the abrupt shift in iOS programming from Objective-C to Swift.
+Fig.
+3c visualizes Python’s growth, highlighting all skills for which it ranks among the top three languages.
+Initially focused on data science, Python spread between 2014 and 2018 into software development areas that range from app development to web-related work, DevOps, and cloud computing.
+Today, it is the dominant language for at least one skill in every cluster.
+As detailed in the Appendix E, this pattern is preserved when we re-weight languages by their GitHub script counts — a proxy for real-world language usage — indicating that Python’s reach extends beyond Stack Overflow.
+4 User expertise
+We can study the expertise of SO users, by analyzing the 237-dimensional skill vectors of eq.
+(1), which count how many answers a user provides to questions associated with each skill in our taxonomy.
+These vectors therewith capture the intensity of engagement in different areas of software work.
+4.1 Voting
+To study whether engagement is related to users’ proficiency in a skill, we analyze how many votes their answers receive.
 Votes are social feedback from other users indicating appreciation for a post.
-However, the sample of answers that users post is highly selected.
-First, before deciding to post an answer, users will look at existing answers to a question to see if they can improve or add to them.
-To avoid sample selection bias from this, we limit ourselves to the first recorded answers for each question.
-Second, users will post answers to question for which they perceive themselves to be sufficiently knowledgeable.
-We address this type of sample selection bias in Appendix D.2, where we use variation in the user base’s activity patterns across the 24-hour daily cycle.
-To determine the importance of users’ task experience, we estimate its effect on the upvotes their answers receives.
-To do so, we estimate the following regression model, with outcomes reported in Fig.
-3c:
+However, before deciding to post an answer, users will first look at existing answers to see if they can improve or add to them.
+As a consequence, the sample of answers that users post is highly selected.
+To avoid that this introduces sample selection bias, we limit ourselves to the first recorded answers for each question.
+Furthermore, users will post answers to question for which they perceive themselves to be sufficiently knowledgeable.
+We address this type of sample selection bias in Appendix D.2, where we exploit exogenous variation in activity patterns throughout the 24-hour daily cycle.
+To assess the effect on the upvotes a user’s answers receive, we estimate the following regression model:
 $y_{a,\theta}=\beta_{x}\log(X_{\theta,u(a)}+1)+\beta_{a}\log(A_{q(a)})+\beta_{v}\log(1+\sum_{i\in Q_{q(a)}}v_{i})+\eta_{a},$ (7)
-where $X_{\theta,u(a)}$ is user $u(a)$’s experience in terms of the number of answers provided to questions involving task $\theta$ in the two calendar years preceding the calendar year in which answer $a$ was provided, $A_{q(a)}$ the number of answers provided to question $q(a)$, $Q_{a}$ the set of answers provided to question $q(a)$, $v_{i}$ the number of upvotes that answer $i$ received and $\eta_{a}$ a disturbance term.
-Furthermore, the dependent variable $y_{a,\theta}$ is either a binary variable that indicates whether answer $a$ is a question’s top answer or $\log(v_{a}+1)$, the logarithm of the number of upvotes answer $a$ receives, augmented by 1 to avoid $\log(0)$ issues.
-Note that if an answer is associated with multiple tasks, we replicate the observation accordingly.
-4.7.3 Task adoption analysis
-To study which new tasks users adopt, we estimate the effect of task value and density on new task entries, estimating the following regression model:
-$y_{u,\theta}=\beta_{V}\log(V_{\theta})+\beta_{D}D_{u,\theta}+\sigma_{u,\theta},$ (9)
-where $y_{u,\theta}$ is a binary variable that encodes whether or not user $u$ adopted task $\theta$, $\log(V_{\theta})$ is the logarithm of the task’s value as defined in eq.( 3), and $D_{u,\theta}$ the user $u$’s density around task $\theta$ as defined in eq.
-(6) before.
-We estimate this model once with data from all programming languages and once with data on answers related to Python only.
+where $X_{\theta,u(a)}$ is user $u(a)$’s expertise in terms of the number of answers provided to questions involving skill $\theta$ in the two calendar years preceding the calendar year in which answer $a$ was provided.
+The next two variables serve as control variables for the general amount of interest a question generates on SO: the first, $A_{q(a)}$, is the number of answers provided to question $q(a)$.
+The second term counts the number of upvotes across all answers to the question, where $Q_{a}$ is the set of answers provided to question $q(a)$ and $v_{i}$ the number of upvotes that answer $i$ received.
+$\eta_{a}$ is a disturbance term.
+Finally, the dependent variable $y_{a,\theta}$ is either a binary variable that indicates whether answer $a$ is a question’s top answer or a continuous variable capturing an answer’s popularity: $\log(v_{a}+1)$, the logarithm of the number of upvotes answer $a$ receives, augmented by 1 to avoid $\log(0)$ issues.
+In some cases, an answer is associated with multiple skills.
+Whenever this happens, we duplicate the observation accordingly, such that this answer generates one observation for each skill involved.
+We expect that users are particularly apt at answering questions that strongly connect to their prior expertise.
+Fig.
+4a supports this conjecture, showing that there is a statistically significant, positive link between users’ skill expertise and the popularity of their answers.
+The estimates plotted in the figure indicate that a 10% rise in skill expertise boosts the likelihood of an answer becoming the top-voted solution by 0.1 pp ($\pm 0.004$) and increases the number of votes it receives by 0.14% ($\pm 0.006\%$).
+In the Appendix D.2, we show that correcting for sample-selection issues increases estimated effects to 0.6 pp ($\pm 0.35$) and 0.9% ($\pm 0.41\%$), respectively.
+Together, these results corroborate the value of skill vectors as summaries of a user’s expertise.
+4.2 Skill acquisition
+To study how user learn new skills, we randomly divide the population of users into two equally sized samples.
+We use the first sample, $U_{1}$, to construct a skill relatedness matrix from co-occurrence patterns of skills in users.
+The second sample, $U_{2}$, is used to estimate how relatedness affects user activity on SO.
+To do so, we collect information from all answer posts by these users within rolling two-year intervals.
+That is, we construct expertise vectors $\vec{S}_{u}^{t}$, where $t$ denotes an interval that stretches across the two calendar years preceding year $t$.
+For instance, $\vec{S}_{u}^{2018}$ refers to the answers user $u$ provides between January 1, 2016 and December 31, 2017.
+We collect these vectors year-by-year across all users in sample $U_{2}$ in matrices $T^{t}$.
+Analogously to eq. (5), we calculate for all users, $u\in U_{2}$, and years, $t$, their expertise density around each skill, collecting results in user-skill matrix $D^{t}$:
+$\mathrm{D}^{t}=\bar{\mathrm{R}}^{U_{1}}\mathrm{T}^{t},$ (8)
+where $\bar{\mathrm{R}}^{U_{1}}$ is the row-normalized relatedness matrix, constructed from the skill vectors of users in sample $U_{1}$.
+Next, we analyze how SO users develop new expertise over time.
+To do so, we first test whether the relatedness among skills—reflected in the skill space—predicts which new skills a user will acquire.
+We take all skills in which user $u$ showed no activity in the two calendar years preceding $t$.
+These skills represent the set of skills that user $u$ has not yet acquired.
+We tag all such skills with a $1$ whenever the user posts at least one answer related the the skill in year $t$, and with 0 otherwise.
+Next, we divide all user-skill combinations where a skill is at risk of being acquired by the user into 10 equally sized bins based on their density values.
+We then estimate the probability that users acquire a skill in year $t$ as the relative frequency of 1s in each bin.
+Standard errors are calculated analogously to eq.
+(6).
+Fig.
+4b shows how these estimated probabilities rise with the user’s expertise density around a skill.
+Users are most likely to pick up skills closely related to those they already master, with the probability of acquiring a new skill rising from 0.9% (when the skill is unrelated) to 14.3% (when it is highly related).
+Apart from a close alignment with their existing skills, higher rewards may also matter.
+We therefore next regress a dichotomous variable that encodes skill acquisition on the skill value estimated in eq.
+(3) and the user’s density around the skill:
+$y_{u,\theta,t}=\beta_{V}\log(V_{\theta})+\beta_{D}D_{u,\theta,t}+\sigma_{u,\theta,t},$ (9)
+where $y_{u,\theta,t}$ is a binary variable that encodes whether or not user $u$ acquires skill $\theta$ in year $t$, $\log(V_{\theta})$ the logarithm of the skill’s value, and $D_{u,\theta,t}$ the user $u$’s density around skill $\theta$ in the two years preceding year $t$, as defined in eq.
+(8).
+We estimate this model once with data from all programming languages and once with data that rely only on answer posts related to Python.
+Figure 4: a.
+Regression analysis of answer popularity on skill expertise.
+Popularity is measured as the number of votes the answer receives, $\log(\#\text{votes}_{a}+1)$ or whether or not the answer is the top answer to the question, skill expertise as the number of prior answers the user provided to questions requiring skill $\theta$ in the preceding two years, $\log(\#\text{answers}_{u(a),\theta,t(a)})$.
+Control variables include $\log(\#\text{answers}_{q(a)})$ and $\log\left(\sum_{\alpha\in A_{q}}\#\text{votes}_{\alpha}+1\right)$, the total number of answers provided to question $q(a)$, and the sum of all votes across these answers.
+To avoid problems due to $\log 0$ values, we add 1 to counts that can evaluate to 0.
+The plot shows point estimates with their 95% confidence intervals.
+Orange markers refer regression analyses of whether or not an answer is the top answer, purple markers to analyses of the number of votes.
+b.
+Estimated probability of acquiring new skills at different values of density, users’ relatedness-weighted expertise in other skills: $d_{\theta,u}=\sum_{\kappa}\frac{\mathrm{R}_{\theta,\kappa}}{\sum_{\tau}\mathrm{R}_{\theta,\tau}}X_{\kappa,u}$, where $X_{\kappa,u}$ denotes user $u$’s prior expertise in skill $\kappa$ and $\mathrm{R}_{\theta,\kappa}$ the relatedness between skills $\theta$ and $\kappa$.
+c.
+Effect of skill value on the probability that a user acquires a skill from a regression analysis of whether or not a user will acquire a new skill $\theta$, on the skill’s value, $V_{\theta}$, controlling for the user’s density around the skill, $d_{\theta,u}$.
+Orange markers refer to analyses that use Python-related questions only, purple markers to the full sample.
 Findings are summarized in Fig.
-3d, with detailed results described in Appendix D.3.
-Acknowledgements
-XF, JW, SD and FN acknowledge financial support from the Austrian Research Agency (FFG), project #873927 (ESSENCSE).
-JW acknowledges support from the Hungarian National Scientific Fund (OTKA FK 145960).
-SD acknowledges financial support from the Dutch Research Council (NWO) under the REWIRE project, grant number KICH1.MV02.22.010.
+4c, with detailed results described in Appendix D.3.
+Somewhat surprisingly, developers do not universally pursue high-value skills.
+As shown in the figure’s right-most panel, this result holds even when we control for how related these skills are to the user’s current skill sets.
+This may reflect higher technological barriers or learning costs associated with high-value skills.
+Interestingly, however, Python users follow a different pattern: when we restrict the analysis to Python-related questions (shown by purple markers), skill value has a positive effect on their decision to learn a new skill.
+This suggests that Python’s versatility allows its users to more readily acquire profitable expertise.
+This not only offers an explanation for Python’s popularity, but also highlights its distinct role in the software ecosystem: as a general-purpose language, it enables users to branch into new, higher-value expertise areas more easily, altering and expanding their career trajectories.
+5 Discussion
+The growing complexity of the labor market has made measuring worker capabilities both more difficult and more important.
+Here we demonstrated that data on individual problem solving can be used to generate a useful, fine-grained taxonomy of skills used in software-related work, a key industry in the digital economy.
+These skills can be given coherent economic values that help predict wages in job advertisements and be used to guess which questions users provide the best answers to according to feedback from the community.
+This illustrates the validity and usefulness of the skill taxonomy inside and outside the context of SO.
+The fine-grained detail of the skill taxonomy allows us to assess the mapping of programming languages to skills.
+This reveals a nested landscape of skills, where ubiquitous skills are undemanding in terms of which programming language they require, but specialized skills require not custom-tailored languages but are rather implemented in generalist, highly versatile, languages.
+A core insight from our results is that skills exist in a space of related capabilities, echoing network-based accounts of human capital (Anderson, 2017; Alabdulkareem et al., 2018; Stephany and Teutloff, 2024).
+Constructing a network based on SO postings that connects skills that are often possessed by the same users shows that real-world software development jobs are coherent, requiring skills that are closely related to one another.
+Moreover, the topology of this network steers how programmers learn: over time SO users acquire new skills that are closely related to the ones they already possess.
+Another important predictor of skill acquisition is a skill’s value.
+Somewhat surprisingly, in general, SO users branch into new skills that are of relatively low value.
+Presumably, these skills are easier to learn, which may explain their modest returns in the first place.
+However, this behavior is different for Python users, who preferentially acquire high-value skills.
+Our networked taxonomy of software skills thereby provided a micro-to-macro explanation for the rise of the Python programming language: Python enables users to learn more valuable skills.
+This interpretation is consistent with broader indicators of Python’s role in the ecosystem.
+Python ranks among the most widely used languages in major indices (TIOBE Software BV, 2026), and developer surveys emphasize that it is used across many application areas (JetBrains and Python Software Foundation, 2024).
+In addition, many AI-assisted data analysis tools solve tasks by executing Python code in a sandboxed environment (OpenAI, ), which makes Python a natural interface between developers and expanding AI-related workflows.
+Empirical studies have shown that Python is one of the most accessible and ubiquitous languages (Juhász et al., 2026).
+Taken together, these patterns fit an account in which Python lowers the cost of moving into high-value skill clusters.
+Our findings connect with the literature on skill complexity, labor market transitions, and the organization of work, complementing studies that have emphasized how granular human capital analysis can reveal hidden dynamics in career mobility, wage disparities, and skill complementarities (Autor et al., 2003; Anderson, 2017; Alabdulkareem et al., 2018; del Rio-Chanona et al., 2021; Stephany and Teutloff, 2024).
+Previous research however, tends to use standardized classifications such as O*NET to define skills and applies them to curated job descriptions, limiting the granularity and responsiveness of skill taxonomies.
+In contrast, by extracting skills directly from real-world problem-solving interactions on SO, our approach captures the immediate and practical realities of skill application at a global scale.
+The value of this approach is demonstrated in its ability to study and explain phenomena from the micro level (the coherence of job requirements and fine-grained individual learning trajectories) to the macro level (the mapping of skills to programming language and the rapid rise to dominance of Python) linking individual to collective learning dynamics.
+We highlight three possible applications of our method and taxonomy.
+First we have demonstrated that microdata on activity can be fruitfully aggregated into skills to describe skill acquisition, development and transitions.
+Skills from any sector in which large-scale fine-grained information on individual behavior at work is available could be categorized in a similar manner.
+Second, in the software sector, employers can use our approach to identify skills related to work activities their employees actually need to perform, enabling better labor market matching outcomes and internal skill development programs.
+Finally, educators can better understand the labor market landscape of the software industry using our taxonomy.
+There are several limitations to our study, which may be tackled in future work.
+First, we have focused exclusively on programming-related skills.
+However, software developers also need many other skills, such as management and reporting capabilities.
+Although SO contains some postings about organizational aspects of software development, its emphasis is on technical subjects.
+To explore how programming skills are combined with other types of expertise in software jobs, future studies could use job ads to compare the co-occurrence of our programming skills with different kinds of soft skills.
+Second, although the question-answer structure makes SO ideal for identifying software work activities and their prerequisite skills, it may be less suited for descriptions of macro-level trends and individual career paths.
+In Appendix E, we compare user numbers on SO to GitHub.
+This analysis suggests that in terms of capturing software developers, both datasets are closely aligned.
+In this Appendix, we therefore also rescale the activity in different programming languages such that they reflect their presence on GitHub.
+However, it is impossible to judge the representativity of SO at the level of individual programming skills.
+Moreover, recently, SO usage has strongly declined due to introduction of generative AI and advanced coding assistants (del Rio-Chanona et al., 2024).
+In future work, we therefore plan to identify software skills directly in code posted on GitHub.
+Third, we opted to create a skill taxonomy that is time-invariant.
+However, an analysis of how programming skills change would provide a more vivid picture of the software industry.
+Such an analysis could also expand the size of our job ad dataset.
+Although sufficient for the validation exercise we conducted, a larger dataset would allow us to see how skills in software development and their value change over time.
+Notwithstanding these limitations, our study demonstrates how question-and-answer data can inform us about the evolution of software development work by quantifying the abstract notion of a skill.
+Programming skills are rapidly diffusing beyond pure software development firms and understanding these skills and their evolution is pivotal to understanding the future of work in an increasingly digital economy.
+Our application contributes to such understanding by shedding light on novel aspects of this at high granularity, at scale, and in near real-time.
+Data Availability
+Declaration of generative AI use
+The authors acknowledge the use of the following AI tools: (a) ChatGPT-3.5 to extract job requirements and salary information from job ads.
+(b) ChatGPT-4.0 to provide descriptive labels for the skills in our skill taxonomy.
+These skill labels were used to allow for easy reference to skill examples in tables and in the main text and to link the job requirements extracted from job ads as detailed in point (a) above to our skill taxonomy.
+In the latter case, findings were corroborated using links that do not rely on the AI-generated skill labels.
+All prompts are provided in the main content.
 Appendix A Stochastic block models
@@
 These models typically postulate a stochastic data generating process of the network that produces communities and the edges among nodes within and between these communities.
-Under this assumption, the observed network can be described as sampled from this process.
-More precisely, given an $N$-nodes network, $G$, and partition or community structure ${\bf c}$ assigning nodes to different groups, the probability of generating the graph $G$ from its partition ${\bf c}$ is:
+Observed networks are assumed to be sampled from this process.
+More precisely, given an $N$-node network, $G$, and partition or community structure, $\bf c$} that assigns nodes to different groups, the probability of generating graph $G$ from its partition $\bf c$} is:
 $P(G|\bf c,\theta),$ (10)
 where $\theta$ is a set of parameters that govern how nodes connect to one another, depending on the communities they belong to.
-The inference procedure inverses this relation, such that it recovers the probability that a specific generative model is correct, given the observed network $G$.
-Thus, the probability of observing the configuration given observed network $G$ with model parameters that observe prior probabilities $P(\theta,{\bf c})$ is:
+Inference inverses this relation, such that it recovers the probability that a specific generative model is correct, given the observed network $G$.
+Thus, the probability of observing the configuration, given the observed network $G$ with model parameters that observe prior probabilities $P(\theta,{\bf c})$ is:
 $P({\bf c}|G)=\frac{\sum_{\theta}P(G|\theta,{\bf c})P(\theta,{\bf c})}{P(G)},$ (11)
-Under the assumption that the generative model has only one set of particular parameters for network $G$, the equation simplifies to:
+Under the assumption that the generative model has only one set of parameters for network $G$, the equation simplifies to:
 $P({\bf c}|G)=\frac{P(G|\theta,{\bf c})P(\theta,{\bf c})}{P(G)}.$ (12)
@@
 Applying this bipartite SBM to our question-tag network yields communities for both questions and tags.
-In this research, we are only interested in communities of tags (as opposed to questions) of which we discover 247, containing a total of 5,083 tags.
+In the current paper, we are only interested in communities of tags (as opposed to questions) of which we discover 247, containing a total of 5,083 tags.
 The procedure yields a community label for each tag, even when there is some ambiguity over which community fits the tag best.
@@
 To reduce this noise, we try to remove tags that are only weakly connected to the other tags in their community.
-In principle, SBM models can identify such nodes using MCMC simulations.
-However, the size of our question-tag renders this approach computationally infeasible.
+In principle, SBMs can identify such nodes using MCMC simulations.
+However, the size of our question-tag matrix renders this approach computationally infeasible.
 Therefore, we here use a simpler approach.
@@
 Next, we label nodes in this tag network according to the community structure we discovered above.
-For each tag, we now calculate the overrepresentation, $O_{tc}$ of (weighted) links to nodes in each of the 247 communities:
+For each tag, we now calculate the overrepresentation, $O_{tc}$, of (weighted) links to nodes in each of the 247 communities:
 $O_{tc}=\frac{w_{tc}/\sum_{c^{\prime}}w_{tc^{\prime}}}{\sum_{t^{\prime}}w_{t^{\prime}c}/\sum_{t^{\prime},c^{\prime}}w_{t^{\prime}c^{\prime}}},$ (13)
@@
 To focus on core tags, for each community, we remove the $20\%$ of tags with the lowest $O_{ct}$ values.
-Finally, we also remove communities with fewer than 3 tags, for which the tag composition provides too little information on the associated task.
-In this way, we arrive at a task space with 237 tasks, composed of a total of 4,054 tags.
-Appendix B Task space
-In this section, we provide a detailed, annotated version of the task space depicted in Fig. 1b. This annotated task space is shown in Fig. B.1.
-Each node is labeled with an example of a programming task provided by ChatGPT-4.0. To do so, we prompt the LLM as follows:
+Finally, we also remove communities with fewer than 3 tags, for which the tag composition provides too little information on the associated skill.
+In this way, we arrive at a skill space with 237 skills, composed of a total of 4,054 tags.
+Appendix B Skill space
+In this section, we provide a detailed, annotated version of the skill space depicted in Fig. 1b. This annotated skill space is shown in Fig. B.1.
+Each node is labeled with an example of a programming skill provided by ChatGPT-4.0. To do so, we prompt the LLM as follows:
 I have a list of tags related to software engineering and computer science.
-The tags are ordered by their importance, listing the most important task first.
+The tags are ordered by their importance, listing the most important tags first.
 Using no more than 10 words, could you please suggest a general task related to the following tags ’task_description’ as key words?
@@
 Please only return json content and not anything else.
-Figure B.1: Fully annotated task space.
+Because questions describe hands-on coding problems, this prompt refers to tasks (generalized problems), which we regard as closely related to the skills they require.
+Doing so avoids that the LLM tries to guess which skills are required to solve the problem stated in a question.
+Figure B.1: Fully annotated skill space.
+Fig.
+B.2a illustrates how SO activity shifted across software expertise areas between 2009 and 2022.
+Notable growth appears in Web Frameworks, certain Android-related skills, Cloud Computing, and AI/ML.
+Conversely, skills associated with Advanced Programming Concepts (e.g., skills to Develop a data structure library for efficient data manipulation, Develop a multimedia application in Windows, or Develop a thread-safe application to prevent race conditions), as well as to Networking or Enterprise and Web App Development, have seen their relative importance on SO decline.
+Figure B.2: Skill dynamics.
+a.
+Skill user-share change from 2009 to 2022.
+Orange markers signal increases, purple markers decreases, in user-shares between 2009 and 2022.
+Marker transparency reflects the size of shifts: darker tones indicate larger changes in user shares.
 Appendix C Analyzing job advertisements
@@
 The analyses in Fig.
-2 of the main text, we assign SO tasks to job advertisement.
-To do so, we first feed the job ad’s text to a large language model (ChatGPT-3.5) to extract strings that describe all task requirements from the listed job description.
-Next, we use sBert[Reimers and Gurevych, 2019] to embed each of the job ad’s task requirements in an 384 dimensional space.
-We then do the same for SO task labels we generated in Sec.
+2 of the main text, we assign skills to job advertisement.
+To do so, we first feed the job ad’s text to a large language model (ChatGPT-3.5) to extract strings that describe all work activities (“tasks”) and salary information from the listed job description, with the following prompt:
+Here is a job ad:{text}. Could you extract in a json format the following information with required json key names:
+1. company name, with key name ‘company_name’.
+2. bool of whether the ad is for more than one job, with key name ’multi_jobs’.
+3.
+With key name ’jobs’, build a list containing the information of each jobs: for each job, build a key-values pair with the job name as key name and value containing the following items: bool of whether this ad contain information of yearly salary values, with key name ’salary_index’; location, with key name ’location’; salary, with key name ’salary’ and containing the following items: if the salary is a range, please give an average value with the key name ’salary_average’, the minimum value with the key name ’salary_min’, the maximum value with the key name ’salary_max’; if the salary is not a range, please give the value with the key name ’salary_average’; currency of salary, with key name ’currency’; bool of whether salary includes equity, with key name ’include_equity’; a list of skills required by the job, with key name ’skills’.
+Please only use the content in the advertisement and do not use other outside information.
+Next, we use sBert[Reimers and Gurevych, 2019] to embed each of the listed work activities in a 384 dimensional space.
+We then do the same for the skill labels we generated in Sec.
 B.
 To increase the accuracy of these embeddings we prompt ChatGPT-4.0 to produce somewhat longer labels of up to 30 words.
-Finally for each task requirement, we pick the closest SO task in terms of the cosine similarity between the two embeddings.
-We keep only those task requirements for which the matched SO task’s cosine similarity exceeds 0.3.
-Below, we illustrate this using a concrete example of a job advertisement. The job advertisement contains the following text:
+Finally for each work activity, we pick the closest skill in our taxonomy in terms of the cosine similarity between the two embeddings.
+We keep only those work activities for which the matched skill’s cosine similarity exceeds 0.3.
+Below, we illustrate this using an example of a concrete job advertisement. The job advertisement contains the following text:
 XXX is a home security and automation company that combines industry-leading technology and truly personal service to protect our customers.
@@
 The things we work on are clearly tied to goals, and …
-The task requirements the LLM extracts, together with the SO tasks we match to them, are listed in Table C.1.
+The job requirements the LLM extracts, together with the skills we match to them, are listed in Table C.1.
 The procedure leads to plausible matches.
-The first job requirement, AWS (Amazon Web Services, a cloud computing service), is clearly closely related to the matched SO task, which contains several tags related to scheduling tasks, such as job-scheduling or quartz-scheduler and several amazon-specific tags such as aws-application-load-balancer and amazon-athena.
+The first job requirement, AWS (Amazon Web Services, a cloud computing service), is clearly closely related to the matched skill, which contains several tags related to scheduling tasks, such as job-scheduling or quartz-scheduler and several amazon-specific tags such as aws-application-load-balancer and amazon-athena.
 Also for the second job requirement, the identified match is plausible.
 Ansible and Terraform are two platforms for CI/CD (Continuous Integration and Continuous Delivery/Deployment) pipelines, which help automate the building, testing, and deployment of code.
-In this context, Maven is primarily used in Java and Kubernetes focuses on automating the deployment of containerized applications.
-For the third and fourth job requirements, the similarity to their matched SO tasks falls below the threshold value of 0.3.
-Yet, the job requirements and matched SO tasks are still related, albeit somewhat more weakly.
+In this context, Maven, which is primarily used in Java and Kubernetes, focuses on automating the deployment of containerized applications.
+For the third and fourth job requirements, the similarity to their matched skills falls below the threshold value of 0.3.
+Yet, the job requirements and matched skills are still related, albeit somewhat more weakly.
 In the third job requirement, metric aggregation systems are often used to monitor CPU or network loads, which support scalable real-time data processing.
-Similary, the PCI (Payment Card Industry compliance mentioned in the fourth job requirement covers security standards required when handling debit or credit card information.
-This typically involves the implementation of secure user authentication and authorization processes of the matched SO task.
-Table C.1: Tasks requirements in job ads and their matching SO tasks.
-(1) task requirements (2) matched SO task similarity (cosine)
+Similarly, the PCI (Payment Card Industry) compliance mentioned in the fourth job requirement covers security standards required when handling debit or credit card information.
+This typically involves the expertise in implementing secure user authentication and authorization processes related to the matched skill.
+Table C.1: Matching job requirements to skills.
+(1) job requirements (2) matched skill similarity (cosine)
 AWS Deploy scalable cloud-based application using AWS services 0.7987
@@
 PCI compliance Implement secure authentication and authorization in applications 0.2487
-Column (1): required tasks as extracted by the LLM from the job ad. Column (2): the label of the matched SO tasks. Column (3): cosine similarity between the embedding vectors of Columns (1) and (2).
-C.2 Salary versus task value
-To assess how well SO tasks predict salaries in job ads, we focus on the subset of job advertisements that also provide salary information.
-For these jobs, we calculate the total number of SO tasks they list, as well as the average value of these tasks.
+Column (1): job requirements as extracted by the LLM from the job ad. Column (2): the label of the matching skills. Column (3): cosine similarity between the embedding vectors of Columns (1) and (2).
+C.2 Salary versus skill value
+To assess how well our skill taxonomy can be used to predict salaries in job ads, we focus on the subset of job advertisements that also provide salary information.
+For these jobs, we calculate the total number of skills we match to them, as well as the average value of these skills.
 Next, we estimate the following regression model:
 $\log(w_{j})=\beta_{v}\bar{v}_{j}+\beta_{r}\bar{r}_{j}+\beta_{n}\log(n_{j})+\tau_{y(j)}+\varepsilon_{j},$ (14)
-where $w_{j}$ is the salary or midpoint of the salary range posted for job $j$, $\bar{v}_{j}$ is the average of the logarithm of the task values across all SO tasks job $j$ requires, $\bar{r}_{j}$ the average of the task density around these SO tasks (i.e., the coherence of the tasks the job requires) and $n_{j}$ the number of SO tasks job $j$ requires.
+where $w_{j}$ is the salary or midpoint of the salary range posted for job $j$, $\bar{v}_{j}$ is the average of the logarithm of the skill values across all skills that job $j$ requires, $\bar{r}_{j}$ the average of the skill density around these skills (i.e., the coherence of the skills the job requires) and $n_{j}$ the number of skills that job $j$ requires.
 $\tau_{y(j)}$ represents year fixed effects to control for general trends in the labor market.
@@
 Outcomes are reported in the table of Table C.2.
-The estimated effect effect of task value on posted salaries is around 0.9 across all models, suggesting that a $10\%$ increase in task value is associated with a $9\%$ higher salary.
+The estimated effect effect of skill value on posted salaries is around 0.9 across all models, suggesting that a $10\%$ increase in skill value is associated with a $9\%$ higher salary.
 Table C.2: Salary regression in online job ads.
@@
 Variables
-task value 0.8734∗∗∗ 0.8522∗∗∗ 0.8853∗∗∗ 0.8669∗∗∗
+skill value 0.8734∗∗∗ 0.8522∗∗∗ 0.8853∗∗∗ 0.8669∗∗∗
 (0.1594) (0.1550) (0.1591) (0.1550)
-task coherence 0.0775∗∗ 0.0273 0.0253
+skill coherence 0.0775∗∗ 0.0273 0.0253
 (0.0294) (0.0257) (0.0262)
-# tasks -0.0059 -0.0067 -0.0066
+# skills -0.0059 -0.0067 -0.0066
 (0.0034) (0.0032) (0.0033)
@@
 p-value: ***: 0.01, **: 0.05, *: 0.1.
-The dependent variable is the value or midpoint of the salary range posted in a job ad, task value is the average of the logarithm of the task values across all SO tasks the job requires, task coherence is the average of the task density round all SO tasks the job requires, # tasks is the logarithm of the number of SO tasks the job requires.
-C.3 Alternative job task matches and thresholds
-We test three alternative methods to match task requirements in job ads to SO tasks.
-To do so, let $v_{t}^{j}$ denote the embedding vector of task requirement $t$ of job $j$.
-Our first method, used in the main text, calculates the cosine similarity between $v_{t}^{j}$ and the embedding vector of each of the SO task’s (long) labels, and then selects the SO task that yields the highest cosine similarity.
-Our second method uses the similarity of $v_{t}^{j}$ to the embedding of the most prominent tag (in terms of its usage on SO) in each SO task.
-Our third method calculates the similarity of $v_{t}^{j}$ to the average embeddings of the tags defining the SO task.
-Finally, our fourth method uses the similarity of $v_{t}^{j}$ to the embedding of the closest tag in each SO task.
+The dependent variable is the value or midpoint of the salary range posted in a job ad, skill value is the average of the logarithm of the skill values across all skills the job requires, skill coherence is the average of the skill density round all skills the job requires, # skills is the logarithm of the number of skills the job requires.
+C.3 Alternative skill matches and thresholds
+We test three alternative methods to match job requirements in job ads to skills.
+To do so, let $v_{t}^{j}$ denote the embedding vector of job requirement $t$ of job $j$.
+Our first method, used in the main text, calculates the cosine similarity between $v_{t}^{j}$ and the embedding vector of each skill’s (long) labels, and then selects the skill that yields the highest cosine similarity.
+Our second method uses the similarity of $v_{t}^{j}$ to the embedding of the most prominent tag (in terms of its usage on SO) in each skill.
+Our third method calculates the similarity of $v_{t}^{j}$ to the average embeddings of the tags defining a skill.
+Finally, our fourth method uses the similarity of $v_{t}^{j}$ to the embedding of the closest tag in each skill.
 Next, we repeat the analysis depicted in Fig.
 2 with each of these methods.
-Figure C.1 presents the results.
-Figure C.1: Alternative methods to match task requirements in job ads to SO tasks.
+Fig.
+C.1 presents the results.
+Figure C.1: Alternative methods to match job requirements in job ads to skills.
 a,c,e,g.
-Prediction of task requirements.
-$20\%$ of task vector elements are masked and grouped into 10 equally sized bins based on the average relatedness of the masked task to the (unmasked) SO task requirements of a job.
-Plot displays the estimated probability, with its $95\%$ confidence interval, that tasks in a bin are required in the job.
+Prediction of skill requirements.
+$40\%$ of skill vector elements are masked and grouped into 10 equally sized bins based on the average relatedness of the masked skill to the (unmasked) skill requirements of a job.
+Plot displays the estimated probability, with its $95\%$ confidence interval, that skills in a bin are required in the job.
 b,d,f,h.
 Predictions of wage offers.
-Jobs are grouped into equally sized bins based on the average value of the tasks they require.
+Jobs are grouped into equally sized bins based on the average value of the skills they require.
 Vertical axis shows the estimated mean salary posted in the job ads, with their $95\%$ confidence intervals.
 a,b.
-Task requirements are matched to SO tasks using the embedding vector of the SO task’s label, namely the results shown in Fig 2.
+Job requirements are matched to skills using the embedding vector of the skill’s label, namely the results shown in Fig 2.
 c,d.
-Task requirements are matched to SO tasks using the embedding vector of the SO task’s main (most frequent) tag.
+Job requirements are matched to skill using the embedding vector of the skill’s main (most frequent) tag.
 e,f.
-Task requirements are matched to SO tasks using the average embedding vector across all of the SO task’s tags.
+Job requirements are matched to skill using the average embedding vector across all of the skill’s tags.
 g,h.
-Task requirements are matched to SO tasks using the closest embedding vector across all tags of SO tasks.
-In the main text, we only keep task requirements that can be matched with SO tasks at a cosine similarity of $0.3$ or above.
+Job requirements are matched to skills using the closest embedding vector across all tags of a skill.
+In the main text, we only keep job requirements that can be matched to skills at a cosine similarity of $0.3$ or above.
 To explore the significance of this threshold, we repeat our analyses with two another thresholds: $0.2$ and $0.4$.
@@
 C.2.
-Varying the threshold around the $0.3$ value of the main text does not meaningfully alter the outcomes of this analysis: density around masked tasks is strongly predictive of jobs demanding these tasks and salaries posted in job ads rise with the average values of the SO tasks we identify for them.
-Figure C.2: Results of task and salary prediction at varying matching thresholds.
+Varying the threshold around the $0.3$ value of the main text does not meaningfully alter the outcomes of this analysis: density around masked skills strongly predicts which other skills a job demands and salaries posted in job ads rise with the average values of the skill we match to them.
+These alternative matching methods and thresholds show that outcomes are robust and do not depend on the LLM-generated skill labels.
+Figure C.2: Results of skill and salary predictions at varying matching thresholds.
 a,c,e.
-Probability that a masked SO task is listed among a job’s task requirements.
-Masked tasks represent $40\%$ of task vector elements and grouped into 10 equally sized bins based on the average relatedness of (density around) the masked task to the (unmasked) task requirements of a job.
+Probability that a masked skill is required in a job.
+Masked skills representing $40\%$ of skill vector elements are grouped into 10 equally sized bins based on the density around the masked skill of unmasked skill requirements in a job.
 Plot displays estimated probabilities in each bin.
 b,d,f Predictions of salaries.
-Jobs are grouped into equally sized bins based on the average value of their required matched SO tasks.
+Jobs are grouped into equally sized bins based on the average value of their required skills.
 Vertical axis shows the estimated means in each bin of the advertised wage offers.
 a,b.
-Task requirements are matched to SO tasks with a minimum cosine similarity of $0.3$, namely the results shown in Fig.
+Job requirements are matched to skills with a minimum cosine similarity of $0.3$, namely the results shown in Fig.
 2.
 c,d.
-Task requirements are matched to SO tasks with a minimum cosine similarity of $0.2$.
+Job requirements are matched to skills with a minimum cosine similarity of $0.2$.
 e,f.
-Task requirements are matched to SO tasks with a minimum cosine similarity of $0.4$.
+Job requirements are matched to skills with a minimum cosine similarity of $0.4$.
 Vertical bars represent 95% confidence intervals.
@@
 In Fig.
-3c of the main text, we study how the popularity of an answer relates to the task experience of the user that provided it.
-The figure shows that answers by users with more task experiences tend to receive more upvotes and are more likely to be the top answer to a question.
+4a of the main text, we study how the popularity of an answer relates to the skill expertise of the user that provided it.
+The figure shows that answers by users with more skill-specific expertise tend to receive more upvotes and are more likely to be the top answer to a question.
 Table D.1 provides the full set of regression results related to this analysis.
@@
 Variables
-task experience 0.0097∗∗∗ 0.0104∗∗∗ 0.0130∗∗∗ 0.0137∗∗∗
+skill expertise 0.0097∗∗∗ 0.0104∗∗∗ 0.0130∗∗∗ 0.0137∗∗∗
 (0.0008) (0.0002) (0.0016) (0.0003)
@@
 (0.0036) (0.0019) (0.0199) (0.0037)
-# votes all answers -$0∗∗$0402∗∗∗ 0.8604∗∗∗ 0.8596∗∗∗
+# votes all answers -0.0400∗∗∗ -0.0402∗∗∗ 0.8604∗∗∗ 0.8596∗∗∗
 (0.0023) (0.0006) (0.0037) (0.0011)
@@
 year Yes Yes
-year-task Yes Yes
+skill-year Yes Yes
 Fit statistics
@@
 The dependent variable is either a binary variable that encodes whether or not an answer has received the most upvotes among all answers provided to a question, or the logarithm of the number of upvotes the answer received.
-task experience: logarithm of task experience, the total number of answers the users provided to questions related to the focal task in the past two calendar years.
+skill expertise: logarithm of skill expertise, the total number of answers the users provided to questions related to the focal skill in the past two calendar years.
 # answers: logarithm of the total number of answers provided to the focal answer’s question.
 # votes all answers: logarithm of the total number of votes received by these answers.
-Regression models contain either year or task-year fixed effects.
+Regression models contain either year or skill-year fixed effects.
 Because for questions that receive only a single answer, this answer is automatically the top answer, we repeat the analysis in a sample that excludes any question that receives only one answer.
-Table D.2 shows that this leads to larger observed effects of task experience.
+Table D.2 shows that this leads to larger observed effects of skill expertise.
 Table D.2: Voting regressions when questions receive at least two answers.
@@
 Variables
-task experience 0.0260∗∗∗ 0.0279∗∗∗ 0.0313∗∗∗ 0.0336∗∗∗
+skill expertise 0.0260∗∗∗ 0.0279∗∗∗ 0.0313∗∗∗ 0.0336∗∗∗
 (0.0009) (0.0004) (0.0026) (0.0007)
@@
 year Yes Yes
-task-year Yes Yes
+skill-year Yes Yes
 Fit statistics
@@
 As a consequence, the answers we observe on SO are no random sample of user-answer combinations.
-For instance, we expect that users who believe they have little to add to the existing answers to a given question will tend to refrain from posting their own answers.
+For instance, we expect that users who believe they have little to add to existing answers to a given question will tend to refrain from posting their own answers.
 This bias is addressed by only focusing on the first answer (in a temporal sense) that each question received.
@@
 Instead, answers will be provided by users who believe they have useful insights to the problem at hand.
-To address the sample selection bias this causes, we will exploit the fact that, due to timezone differences, most users are active on the SO platform at somewhat predictable times.
+To address the sample selection bias this causes, we will exploit the fact that, due to timezone differences, most users are active on the SO platform at different, yet somewhat predictable, times.
 Therefore, we expect that who will provide the first answer to a question to some extent depends on the time of day the question was posted.
 This introduces some exogenous variation in the likelihood that a given user provides an answer to a given question.
-Moreover, as far as task experience is not fully equally distributed across timezones, this offers exogenous variation in the task experience associated with the first answer to a question.
+Moreover, as far as skill expertise is unequally distributed across timezones, this offers exogenous variation in the skill expertise associated with the first answer to a question.
 We proceed as follows.
@@
 This offers a coarse indication of when a user is active on the platform.
-Finally, we calculate for each task and each day-minute the amount of task experience that is expected to be available on SO.
-To do so, we sum the task experience of all users, weighted by the gap between the day-minute of a question and the average day-minute of each user.
-We call this variable, $M_{\theta,m}$, the minute-task count of task $\theta$ at day-minute $m$.
-Next, we use this variable as an instrument for the actual task experience of whichever user provides the first answer to the question at hand.
+Finally, we calculate for each skill and each day-minute the amount of skill expertise that is expected to be available on SO.
+To do so, we sum the skill expertise of all users, weighted by the gap between the day-minute of a question and the average day-minute of each user.
+We call this variable, $M_{\theta,m}$, the minute-skill count of skill $\theta$ at day-minute $m$.
+Next, we use this variable as an instrument for the actual skill expertise of whichever user provides the first answer to the question at hand.
 Note that there are some potential concerns about this instrument.
-First, if there are geographic components to task experience of users and if these geographic components directly influence the quality of an answer, the instrument may be invalid.
+First, if there are geographic components to skill expertise of users and if these geographic components directly influence the quality of an answer, the instrument may be invalid.
 For instance, the timezone that includes Silicon Valley may have a disproportionate number of highly skilled programmers, such that any questions posed when most people in this timezone are active may attract high quality first answers.
-To correct for this, we always include day-minute fixed effects and in some specifications also add task-year-combination fixed effects.
-Second, the instrument assumes that questions are answered within day from the time they are posted on SO.
-In fact, we would expect that our instrument does not work for questions whose first answer does not arrive within 24 hours.
+To correct for this, we always include day-minute fixed effects and in some specifications also add skill–year-combination fixed effects.
+Second, the instrument assumes that questions are answered within a day from the time they are posted on SO.
+In fact, we would expect that our instrument does not work for questions whose first answer does not arrive within the first 24 hours after the question was raised.
 We exploit this in placebo tests that focus on questions whose first answer is provided at least 24 hours after the posting of the question itself.
@@
 $y_{a,\theta}=\beta_{x}\log(X_{\theta,u(a)}+1)+\beta_{a}\log(A_{q(a)})+\beta_{v}\log(1+\sum_{i\in Q_{q(a)}}v_{i})+\mu_{q(a)}+\eta_{a},$ (15)
-where $u(a)$ is the user who posts answer $a$, $X_{\theta,u(a)}$ user $u(a)$’s experience in terms of the number of answers provided to questions involving task $\theta$ in the two calendar years preceding the calendar year in which answer $a$ was provided, $A_{q(a)}$ the number of answers that will eventually be provided to question $q(a)$, $Q_{a}$ the set of these answers, $v_{i}$ the number of upvotes that answer $i$ received and $\eta_{a}$ a disturbance term.
+where $u(a)$ is the user who posts answer $a$, $X_{\theta,u(a)}$ user $u(a)$’s skill-specific expertise in terms of the number of answers provided to questions involving skill $\theta$ in the two calendar years preceding the calendar year in which answer $a$ was provided, $A_{q(a)}$ the number of answers that will eventually be provided to question $q(a)$, $Q_{a}$ the set of these answers, $v_{i}$ the number of upvotes that answer $i$ received and $\eta_{a}$ a disturbance term.
 Compared to eq.
@@
 As before, the dependent variable $y_{a,\theta}$ is either a binary variable that indicates whether answer $a$ is the top answer to its question or $\log(v_{a}+1)$, the logarithm of the number of upvotes answer $a$ receives, augmented by 1 to avoid $\log(0)$ issues.
-If an answer is associated with multiple tasks, we replicate the observation accordingly.
+If an answer is associated with multiple skills, we replicate the observation accordingly.
 The results is presented in Table D.3.
-The effect of task experience on answer popularity is significant and positive across all model specifications.
+The effect of skill-specific expertise on answer popularity is significant and positive across all model specifications.
 Moreover, the estimated causal effects arrived at by instrumental variables estimation are substantially larger than the OLS estimates reported in the main text.
@@
 Variables
-task experience 0.0162∗∗∗ 0.0601∗∗∗ 0.0306∗∗∗ 0.0919∗∗∗
+skill expertise 0.0162∗∗∗ 0.0601∗∗∗ 0.0306∗∗∗ 0.0919∗∗∗
 (0.0005) (0.0179) (0.0006) (0.0209)
@@
 QA-abs-minute Yes Yes Yes Yes
-task-year Yes Yes
+skill-year Yes Yes
 Fit statistics
@@
 Within R2 0.27051 0.19626 0.80680 0.76881
-Clustered (group_task_qminute) standard-errors in parentheses
+Clustered (group_skill_qminute) standard-errors in parentheses
 Signif. Codes: ***: 0.01, **: 0.05, *: 0.1
@@
 The dependent variable is either a binary variable that encodes whether or not an answer has received the most upvotes among all answers provided to a question, or the logarithm of the number of upvotes the answer received.
-task experience: logarithm of task experience, the total number of answers the user provided to questions related to the focal task in the past two calendar years.
+skill expertise: logarithm of skill expertise, the total number of answers the user provided to questions related to the focal skill in the past two calendar years.
 # answers: logarithm of the total number of answers provided to the focal answer’s question.
 # votes all answers: logarithm of the total number of votes received by these answers.
-Regression models contain day-minute fixed effects (QA-abs-minute) for the timing of the question and year or task-year fixed effects.
-To overcome sample-selection biases, we instrument task experience with $M_{\theta,m}$, the estimated amount of available task experience at the moment the question was posted, calculated as the weighted sum of task experience across all users, where weights reflect the temporal distance between the day-minute of a user and of the question’s posting.
+Regression models contain day-minute fixed effects (QA-abs-minute) for the timing of the question and year or -year fixed effects.
+To overcome sample-selection biases, we instrument skill expertise with $M_{\theta,m}$, the estimated amount of available skill expertise at the moment the question was posted, calculated as the weighted sum of skill expertise across all users, where weights reflect the temporal distance between the day-minute of a user and of the question’s posting.
 Next, we test this strategy using our placebo tests.
@@
 Results are shown in Table D.4.
-For answers provided within 24 hours, task experience has a strong and statistically significant positive effect on answer popularity, regardless of whether we look at the probability of being the top answer or the number of upvotes an answer receives.
-In contrast, for the placebo test, comprising of answers provided more than 24 hours after the question was posted, the effect of task experience becomes insignificant in three out of four models.
-In the model without task fixed effects, the effect on the number of votes an answer receives turns significantly negative.
-This suggests that the task fixed effects control for some relevant confounders.
+For answers provided within 24 hours, skill expertise has a strong and statistically significant positive effect on answer popularity, regardless of whether we look at the probability of being the top answer or the number of upvotes an answer receives.
+In contrast, for the placebo test, comprising of answers provided more than 24 hours after the question was posted, the effect of skill expertise becomes insignificant in three out of four models.
+In the model without skill fixed effects, the effect on the number of votes an answer receives turns significantly negative.
+This suggests that the skill fixed effects control for some relevant confounders.
 Therefore, our preferred estimates are derived from models that include these fixed effects.
-These models suggest a large and positive effect of task experience on the popularity of answers, corroborating our analysis in the main text.
+These models suggest a large and positive effect of skill expertise on the popularity of answers, corroborating our analysis in the main text.
 Table D.4: Instrumental variable regression: placebo tests.
@@
 Variables
-task experience 0.0183∗∗∗ 0.0620∗∗∗ 0.0314∗∗∗ 0.0947∗∗∗ 0.0009 0.0561 -0.0172∗∗∗ -0.0030
+skill expertise 0.0183∗∗∗ 0.0620∗∗∗ 0.0314∗∗∗ 0.0947∗∗∗ 0.0009 0.0561 -0.0172∗∗∗ -0.0030
 (0.0005) (0.0174) (0.0006) (0.0203) (0.0038) (0.2365) (0.0048) (0.2827)
-# answers -$0∗∗$4528∗∗∗ -0.7750∗∗∗ -0.7347∗∗∗ -0.5116∗∗∗ -0.4920∗∗∗ -0.8835∗∗∗ -0.8773∗∗∗
+# answers -0.4800∗∗∗ -0.4528∗∗∗ -0.7750∗∗∗ -0.7347∗∗∗ -0.5116∗∗∗ -0.4920∗∗∗ -0.8835∗∗∗ -0.8773∗∗∗
 (0.0007) (0.0123) (0.0008) (0.0143) (0.0023) (0.0848) (0.0032) (0.1014)
@@
 QA-abs-minute Yes Yes Yes Yes Yes Yes Yes Yes
-task-year Yes Yes Yes Yes
+skill-year Yes Yes Yes Yes
 Fit statistics
@@
 Within R2 0.26317 0.18891 0.79722 0.75538 0.31531 0.22360 0.86502 0.86413
-Clustered (group_task_qminute) standard-errors in parentheses
+Clustered (group_skill_qminute) standard-errors in parentheses
 p-value: ***: 0.01, **: 0.05, *: 0.1
@@
 The results for the latter sample, depicted in the in columns (5)-(8), represent placebo tests for the instrumental variables estimation.
-D.3 Regression of entry on task value
-To study how task values effect users’ adoption on new tasks, we run the regression described in eq.
+D.3 Regression of skill acquisition on skill value
+To study how skill values effect which new skills users learn, we run the regression described in eq.
 (9) of the main text.
-We compare the association of two variables, task value and task density, with the likelihood that users adopt a certain task.
-We repeat this analysis once using all answers to construct experience and task entry variables, and once using only answers to Python related questions.
+We compare the association of two variables, skill value and skill density, with the likelihood that users acquire a certain new skoill.
+We repeat this analysis once using all answers to construct expertise and skill acquisition variables, and once using only answers to Python related questions.
 The results are shown in Table D.5.
-Table D.5: Task entry regressions
-Dependent Variable: entry (binary, all programming languages) entry (binary, Python)
+Table D.5: Skill acquisition regressions
+Dependent Variable: skill acquisition (binary, all programming languages) skill acquisition (binary, Python)
 Model: (1) (2) (3) (4) (5) (6) (7) (8)
 Variables
-task value -0.0078∗∗∗ -0.0079∗∗∗ -0.0068∗∗∗ -0.0069∗∗∗ 0.0342∗∗∗ 0.0348∗∗∗ 0.0140∗∗∗ 0.0146∗∗∗
+skill value -0.0078∗∗∗ -0.0079∗∗∗ -0.0068∗∗∗ -0.0069∗∗∗ 0.0342∗∗∗ 0.0348∗∗∗ 0.0140∗∗∗ 0.0146∗∗∗
 (0.0008) (0.0008) (0.0006) (0.0006) (0.0011) (0.0011) (0.0009) (0.0011)
@@
 (9) of the main text.
-Observations are user-task combinations for which the user did not answer any questions related to the task in the previous two years.
-The dependent variable is a binary value of whether or not the user enters the task (i.e., answers questions related to this task) in the current period.
-task value: logarithm of the task’s imputed value, task density: the user’s prior experience in other tasks, weighted by the relatedness of these tasks to the current task.
+Observations are user-skill combinations for which the user did not answer any questions related to the skill in the previous two years.
+The dependent variable is a binary variable that encodes whether or not the user acquires the skill (i.e., answers questions related to this skill) in the current period.
+skill value: logarithm of the skill’s imputed value, skill density: the user’s prior expertise in other skills, weighted by the relatedness of these skills to the current skill.
 For ease of interpretation, we subtract the variable’s mean and then divide by its standard deviation.
@@
 Appendix E Rescaling by Github
-Figure 4 shows Python’s progressive rise as the dominant language for an increasing number of programming tasks.
-However, even though SO may provide a useful dataset to define such programming tasks, it is not guaranteed to be representative of programmers across the world.
+Fig.
+3 shows Python’s progressive rise as the dominant language used in an increasing number of programming skills.
+However, even though SO may provide a useful dataset to define such programming skills, it is not guaranteed to be representative of programmers across the world.
 To assess this problem, we compare counts of programmers on SO to comparable ones on GitHub.
@@
 In particular, we collect information on the (self-reported) countries of residence on GitHub and (observed) language use between 2020 and 2022 [GitHub, 2025].
-Figure E.1a compares SO to GitHub user counts at the level of countries, languages and country-language combinations.
+Fig.
+E.1a compares SO to GitHub user counts at the level of countries, languages and country-language combinations.
 This shows that user numbers are highly correlated across the two platforms, with correlations ranging from 0.75 to 0.93.
@@
 b.
-Number of tasks in which a programming language is the top language in terms of SO users reweighted by user counts from GitHub, from 2020 to 2022.
+Number of s in which a programming language is the top language in terms of SO users reweighted by user counts from GitHub, from 2020 to 2022.
 The graph shows the largest 6 languages measured by cumulative SO posts between August 2008 and June 2023.
 Nevertheless, some languages are better represented on SO than others.
-This may affect our analysis of how dominant a language is for a given task.
+This may affect our analysis of how dominant a language is in a given skill.
 In Fig.
@@
 Because we only have GitHub information for the years 2020, 2021 and 2022, the analysis is limited to those years.
-Also with reweighted user counts, Python leads in most tasks, albeit with round 50, in somewhat fewer tasks than what we report in the main text.
-The ordering of the remaining languages is mostly unchanged as well, except for C#, which was somewhat more prominent before rescaling user counts.
+Also with reweighted user counts, Python leads in most skills, albeit with around 50, in somewhat fewer skills than what we report in the main text.
+The ordering of the remaining languages also remains mostly unchanged, except for C#, which was somewhat more prominent before rescaling user counts.
 Moreover, also the most prominent temporal patterns over the 3-year time period are mostly left unchanged.
-Figure E.1b retains the rise of Python and the drops of Java and PHP, as well as the relative stability in these years of JavaScript, validating the results shown in the main text.
+Fig.
+E.1b retains the rise of Python and the drops of Java and PHP, as well as the relative stability in these years of JavaScript, validating the results shown in the main text.
 Appendix F Swift vs Objective-C
-Figure F.1: Number of tasks in which Objective-C has more users on SO than Swift and vice versa. The dashed vertical line marks the year 2014, the year that Swift was released.
+Figure F.1: Number of skills in which Objective-C has more users on SO than Swift and vice versa. The dashed vertical line marks the year 2014, the year that Swift was released.
 Software written for Apple devices was mostly written in Objective-C from the 1980s until the mid 2010s.
@@
 Unlike typical programming language dynamics, which evolve organically from developer communities, Swift’s introduction was a deliberate, top-down intervention by Apple.
-Figure F.1 describes this shift in terms of the tasks where Swift was more popular than Objective-C and vice versa.
-Before 2015, most tasks were handled preferentially in Objective-C.
+Fig.
+F.1 describes this shift in terms of the skills where Swift was more popular than Objective-C and vice versa.
+Before 2015, most skills were deployed preferentially using Objective-C.
 However, since the release of Swift in 2014 this rapidly changes.
-Swift surpasses Objective-C in terms of the number of different tasks it handles around 2015 and all but eclipses Objective-C by 2016.
+Swift surpasses Objective-C in terms of the number of different skills in which it dominates around 2015 and all but eclipses Objective-C by 2016.